### v1

# On Discounted Infinite-Time Mean Field Games

Abstract
In this paper, we study the infinite-time mean field games with discounting, establishing an equilibrium where individual optimal strategies collectively regenerate the mean-field distribution. To solve this problem, we partition all agents into a representative player and the social equilibrium. When the optimal strategy of the representative player shares the same feedback form with the strategy of the social equilibrium, we say the system achieves a Nash equilibrium.
We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations. By employing the elliptic master equation, we provide a representation for the Nash equilibrium.
Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, elliptic master equation.
1 Introduction
The study of mean field games was initiated independently by Lasry-Lions (see [8, 9, 10]) and Huang-Malhamé-Caines [7], which is an analysis of limit models for symmetric weakly interacting $(N+1)-$player differential games. It is noteworthy that current theoretical frameworks are primarily developed for finite-time problems, while infinite-time scenarios remain significantly underdeveloped. We refer the reader to [5] for a comprehensive exposition on this subject.
In this paper, we consider a generalized framework for mean field games, which extends the classical finite-time settings to discounted infinite-time mean field games. Our model differs from the fixed-point problems studied in prior works. In our framework, a representative player interacts with a continuum of other players (also referred to as the population or social equilibrium). The dynamic of the private state of the representative player is given by
$X^{x}_{t}=x+\int_{0}^{t}b(X^{x}_{t},\mu_{s},\beta_{s}^{x})\mathrm{d}t+B_{t}.$ (1.1)
Here $\mu_{t}$ is the population distribution , while $\beta^{x}$ is the optimal strategy of the representative player, derived as the solution to the stochastic optimization problem:
$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right],$ (1.2)
under the constraint
$\begin{cases}\mathrm{d}X_{t}^{x,\beta}=b(X_{t}^{x,\beta},\mu_{t},{\beta}_{t})\mathrm{d}t+\mathrm{d}B_{t},\\ X_{0}^{x,\beta}=x.\end{cases}$ (1.3)
Since the population consists of a multitude of homogeneous individuals, the macroscopic state should satisfy
$X^{\xi}_{t}=X^{x,\beta^{x}}|_{x=\xi},$ (1.4)
where $\mathcal{L}_{\xi}=\mu_{0}$ denotes the initial population distribution. We say the population reaches equilibrium if $\mu_{t}=\mathcal{L}_{X^{\xi}_{t}}.$
This model decouples the micro-level agent from the macro-level societal distribution, enabling interconnected analysis of their evolution. The equilibrium is characterized by two consistency conditions:
  • Individual Rationality: The representative player’s optimal strategy is consistent with the perceived social equilibrium.
  • Macro Consistency: The aggregate distribution generated by all players adopting this strategy must equal the posited social equilibrium.
We will employ the stochastic maximum principle and infinite-time forward-backward stochastic differential equations (FBSDEs) to solve this game-theoretic problem, and derive a representation of the equilibrium strategy via an elliptic-type master equation.
The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [12, 13] in the early 1990s, which is now a typical tool in stochastic optimization problems. We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time forward-backward stochastic differential equations. The foundational work in [14] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, later extended by [1] to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games. In this paper, we partition all agents into a representative player and the social equilibrium, and characterize the equilibrium state through the following system of infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\xi+b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.5)
$\begin{cases}\mathrm{d}X_{t}^{x}=x+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.6)
Here $\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy$ is the generalized Hamiltonian, and $\hat{\alpha}(x,y)$ is the minimizer of $\mathcal{H}(x,\mu,a,y)$ with respect to $a$ when we assume $\mathcal{H}$ is separable in variables $\mu$ and $a$. The solution $X_{t}^{\xi}$ to Equation (1.5) represents the population’s state process, whose law corresponds to the population distribution $\mu_{t}$. And the solution $X_{t}^{x}$ to Equation (1.6) is the state process of the representative agent after solving the optimization problem. Notably, it exhibits the same feedback structure as the population’s state process. To further elucidate the relationship in Equation (1.4 ), we introduce the elliptic-type master equations.
First introduced by Lions in lectures [11], the parabolic-type master equations appeared in the context of the theory of mean field games. This is a time-dependent equation that bears profound connections with finite-time mean field game theory. Essentially, it describes a strategic interaction between a representative player and the collective environment. When the Nash equilibrium exists, the master equation provides a powerful tool to characterize the equilibrium cost and control pattern of this system. We refer the reader to [5, 3, 6] for a comprehensive exposition on the subject. In this paper, we propose the elliptic-type master equations, which explicitly characterizes the feedback forms of both the representative player and the social equilibrium. This equation takes the following form:
$\begin{split}rU(x,\mu)=&H(x,\mu,\partial_{x}U(x,\mu))+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (1.7)
Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are $W_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law. Under the assumption that the master equation (1.7) admits a solution with sufficient regularity, we derive the following representation for Equation (1.5) and (1.6):
$Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (1.8)
$Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (1.9)
If we further assume that $\tilde{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\mu,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any fixed finite time $T$, we have
$X_{t}^{x}|_{x=\xi}=X_{t}^{\xi},\quad t\in[0,T].$ (1.10)
This is precisely the relationship expressed in Equation (1.4).
2 Preliminaries
We will use the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$. Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any measure on $\mathbb{R}$ with finite second moment.
Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$, define the larger filtered probability space by
$\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}.$ (2.1)
Throughout the paper we will use the probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$. However, when we deal with the distribution-dependent master equation, independent copies of random variables or processes are needed. Then we will tacitly use their extensions to the larger space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}},\tilde{\mathbb{F}})$.
Let ${\cal P}\triangleq{\cal P}(\mathbb{R})$ be the set of all probability measures on $\mathbb{R}$ and let ${\cal P}_{p}(p\geq 1)$ denote the set of $\mu\in{\cal P}$ with finite $p$-th moment. For any sub-$\sigma$-field ${\cal G}\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, ${\cal G}$-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ For any $\mu,\nu\in{\cal P}_{p}$, we define the $W_{p}$–Wasserstein distance between them as follows:
$W_{p}(\mu,\nu)\triangleq\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~{}\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}-$valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
[…]
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$. For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
It has a unique solution $(X_{t},Y_{t},Z_{t})\in L^{2}_{K}(0,+\infty,\mathbb{R}^{3})$ under the following assumptions:
Assumption 2.1
(i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
$|B(t,x,y,\mu)$ $-B(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
$\leq\ell(|x-x^{\prime}|+|y-y^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime})).\quad\mathrm{a.s.}$
(ii) There exist constants $0<K<2\kappa$ such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$
$\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (2.5)
$\leq-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
For the detailed proof, we refer the reader to ([1], Theorem 2.1).
We introduce the Wasserstein space and differential calculus on Wasserstein space. For a $W_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$, its $W_{2}$-Wasserstein derivatives[5](also called Lions-derivative), takes the form $\partial_{\mu}U:(\mu,\tilde{x})\in{\cal P}_{2}\times\mathbb{R}\to\mathbb{R}$ and satisfies:
$U({\cal L}_{\xi+\eta})-U(\mu)=\mathbb{E}\big{[}\langle\partial_{\mu}U(\mu,\xi),\eta\rangle\big{]}+o(\|\eta\|_{2}),\ \forall\ \xi\in\mathbb{L}^{2}(\mathcal{F};\mu),\eta\in\mathbb{L}^{2}(\mathcal{F}).$ (2.6)
Let ${\cal C}^{0}({\cal P}_{2})$ denote the set of $W_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$. For ${\cal C}^{1}({\cal P}_{2})$, we mean the space of functions $U\in{\cal C}^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.6). Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on ×$\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on ×$\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$. Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on ×$\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}^{d}\times{\cal P}_{2}\times\mathbb{R}$.
Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5]. For $i=1,2$, let $dX^{i}_{t}\triangleq b^{i}_{t}dt+\sigma^{i}_{t}dB_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}:={\cal L}_{X^{2}_{t}}$. Fix $T>0$ and let all conditions be restricted to the interval [0,T]. Then we have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big{[}\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\hbox{\rm tr$\,$}\big{(}\partial_{xx}U[\sigma_{t}^{1}]^{2}\big{)}\Big{]}(t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big{(}\tilde{\mathbb{E}}_{{\cal F}_{t}}\big{[}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big{]}\Big{)}\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$
3 Infinite-time mean field games with discounting
In this section, we introduce the infinite-time mean field games with discounting. Let $r>0$ represent the time discount factor and $A\subset\mathbb{R}$ be a convex control space. Define $\mathcal{A}\triangleq L^{2}_{r}(0,\infty,A)$ to be the space of all adimissible controls, and $b,f:\mathbb{R}\times{\cal P}_{2}\times A\rightarrow\mathbb{R}$ are two measurable functions.
We consider a population consisting of a continuum of players, where each individual player strategically interacts with the aggregate distribution formed by all other players to minimize their own cost. Let $\mu_{t}$ denote the population distribution and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ denote the initial state with $\mathcal{L}_{\xi}=\mu_{0}$. The state of the representative player with initial value $x$ is given by
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mu_{t},\beta_{s})\mathrm{d}s+B_{t},$ (3.1)
where $\beta\in\mathcal{A}$ is the strategy remains to be determined. The representative player seeks to minimize the cost
$J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (3.2)
Assuming $\beta^{x}\in\mathcal{A}$ minimizes the cost function, the state process of the representative player is $X^{x,\beta^{x}}_{t}$. Since the representative agent can characterize the strategies of all players in the population, we assert that the following fundamental relationship must hold:
$\mu_{t}=\mathcal{L}_{X^{x,\beta^{x}}_{t}|_{x=\xi}}.$ (3.3)
To solve this mean field game problem, we work under the following assumptions:
Assumption 3.1
(i) $b(x,\mu,a)$ is Lipschitz in $(x,\mu,a)$, and $f(x,\mu,a)$ is of at most quadratic growth in $(x,\mu,a)$. There exists a positive constant $\ell$ such that for any $\mu,\mu^{\prime}\in{\cal P}_{2},x\in\mathbb{R},a\in A$,
$|b(x,\mu,a)-b(x,\mu^{\prime},a)|\leq\ell\mathcal{W}_{2}(\mu,\mu^{\prime}).$ (3.4)
(ii) There exists a constant $\lambda>\ell-r/2$ such that for any $a\in A,\mu\in{\cal P}_{2},x,x^{\prime}\in\mathbb{R}$, it holds that
$(x-x^{\prime})\left(b(x,\mu,a)-b(x^{\prime},\mu,a)\right)\leq-\lambda(x-x^{\prime})^{2}.$ (3.5)
These assumptions jointly ensure that: (i) the state process remains confined within the $L^{2}_{r}$ space, and (ii) the cost functional maintains integrability.
Then we partition all agents into two components: (i) the representative player, who dynamically optimizes their strategy, and (ii) the social equilibrium (or mean field), characterizing the macroscopic state shared by the population. For the social equilibrium, its state is governed by the following SDE:
$X_{t}^{\xi,\alpha}=\xi+\int_{0}^{t}b(X_{s}^{\xi,\alpha},\mathcal{L}_{X_{s}^{\xi,\alpha}},\alpha_{s})\mathrm{d}s+B_{t},$ (3.6)
where $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$. We note that, by assumption 3.1, this SDE has a unique strong solution in $L^{2}_{r}$, see ([1], Proposition 2.2) for more details.
The state of the representative player is governed by
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi,\alpha}},\beta_{s})\mathrm{d}s+B_{t}.$ (3.7)
Here, we also require their control $\beta\in\mathcal{A}.$
The representative player seeks to minimize the cost
$J\left(x,\xi;\alpha,\beta\right)=\mathbb{E}\bigg{[}\int_{0}^{+\infty}\mathrm{e}^{-rt}f\big{(}X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi,\alpha}},\beta_{t}\big{)}\mathrm{d}t\bigg{]}.$ (3.8)
We observe that the cost function depends exclusively on the distribution of $X_{t}^{\xi,\alpha}$ and we have
$\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad J(x,\xi^{\prime};\alpha,\beta)=J(x,\xi;\alpha,\beta),\quad~{}\mbox{for all $x,\alpha,\beta$.}$ (3.9)
Therefore, we may define
$J(x,\mu;\alpha,\beta)\triangleq J(x,\xi;\alpha,\beta),\quad\quad\xi\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.10)
Now, for any $(x,\mu)\in\mathbb{R}\times{\cal P}_{2}$ and $\alpha\in\mathcal{A}$, we consider the infimum
$V(x,\mu;\alpha)\triangleq\inf_{\beta\in\mathcal{A}}J(x,\mu;\alpha,\beta).$ (3.11)
Definition 3.2
We say a function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if the closed-loop controls $\alpha_{s}^{\mu_{0}}={\alpha}^{*}(X_{s}^{\xi,\alpha^{\mu_{0}}},\mathcal{L}_{X_{s}^{\xi,\alpha^{\mu_{0}}}})$,$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi,\alpha^{\mu_{0}}}})$ satisfy
$\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\mu_{0};\alpha^{\mu_{0}},\beta).$ (3.12)
When this Nash equilibrium $\alpha^{*}$ exists, we can define the value function with respect to this Nash equilibrium as:
$V(x,\mu_{0})\triangleq J(x,\mu_{0};\alpha^{\mu_{0}},\alpha^{x}).$ (3.13)
Note that our definition of Nash equilibrium differs from that in [1] and earlier finite-time mean field games. In [1], the mean field game is formulated as a fixed-point problem. The population is assumed to be homogeneous, meaning all agents are identical and thus represented by a single representative player. For a given measure flow $(\mu_{t})_{t\geq 0}$ , the representative player wants to minimize
$J^{\mu}(\alpha)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(t,X_{t},\mu_{t},\alpha_{t})\mathrm{d}t\right],$ (3.14)
under the constraint
$\begin{cases}\mathrm{d}X_{t}=b\left(t,X_{t},\mu_{t},{\alpha}_{t}\right)\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (3.15)
Then we require that the law of $X_{t}$ coincides with $\mu_{t}$, which means we need to find a fixed point.
In this paper, we separate a representative player from the population, who only needs to consider their own optimization problem starting from state $x$. The representative player’s state evolution depends on both their own state $X^{x,\beta}$ and the overall population distribution $\mathcal{L}_{X^{\xi,\alpha}}$. Here we would like to emphasize that since there are a large number of players, any change of a representative player doesn’t impact the measure flow $\mathcal{L}_{X^{\xi,\alpha}}$. Under the existence assumption of the Nash equilibrium $\alpha^{*}$ specified in Definition 3.2, the stochastic dynamics of both the population and representative player are characterized by the following stochastic differential equations:
$\begin{cases}X_{t}^{\xi,\alpha^{*}}=\xi+\int_{0}^{t}b(X_{s}^{\xi,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}},{\alpha}^{*}(X_{s}^{\xi,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}}))\mathrm{d}s+B_{t},\\ X_{t}^{x,\alpha^{*}}=x+\int_{0}^{t}b(X_{s}^{x,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}},{\alpha}^{*}(X_{s}^{x,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}}))\mathrm{d}s+B_{t}.\end{cases}$ (3.16)
We further assume that $\tilde{b}(x,\mu)\triangleq b(x,\mu,\alpha^{*}(x,\mu))$ is Lipschitz continuous in $(x,\mu)$. For any fixed finite time $T$, see [4, 2] , we have
$X_{t}^{x,\alpha^{*}}|_{x=\xi}=X_{t}^{\xi,\alpha^{*}},\quad t\in[0,T].$ (3.17)
This implies that every sample point from the initial population follows the same evolutionary dynamics as our hypothesized representative player, which justifies the mathematical validity of using a single representative player to characterize the behavior of all individuals in the population. Moreover, when all agents in the population adopt the same strategy as the representative player, their collective behavior precisely generates the aggregate distribution $\mathcal{L}_{X_{t}^{\xi,\alpha^{*}}}$ derived from our solution. This justifies why we refer to $X_{t}^{\xi,\alpha^{*}}$ as the social equilibrium.
4 Connection with infinite-time McKean-Vlasov FBSDEs
In this section, we employ the maximum principle to solve the optimization problem for the representative player and then use the infinite-time McKean-Vlasov FBSDE to construct the optimal strategy for the representative player such that the controls of the representative player and the social equilibrium share the same feedback form. Our derivation is based on the following key assumptions on $b,f$:
Assumption 4.1
(i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$. Where $b_{0},f_{0}$ are measurable functions on ×$\mathbb{R}\times{\cal P}_{2}$, $b_{1},f_{1}$ are measurable functions on ×$\mathbb{R}\times A$.
(ii) $b,f$ are differentiable with respect to $(x,a)$ and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
(iii) $H_{0}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)$ is convex with respect to $(x,a)$. $\min_{a\in A}H_{0}(x,\mu,a,y)$ has a unique minimizer $\hat{\alpha}(x,y)$ which is Lipschitz continuous in $(x,y)$.
4.1 Pontryagin’s maximum principle
Assuming the state of the social equilibrium $X_{t}^{\xi}$ is given, we consider the optimization problem for the representative player, whose state is given by
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi}},\beta_{s})\mathrm{d}s+B_{t}.$ (4.18)
The cost functional takes the form
$J(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\mathrm{d}t\right],$ (4.19)
and the representative player wants to solve the minimization problem
$\inf_{\beta\in\mathcal{A}}J(\beta).$ (4.20)
Let us formally derive the maximum principle for the infinite-time control problem. Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ Let
$R_{t}=\lim_{\epsilon\to 0}\frac{X_{t}^{x,\beta+\epsilon\gamma}-X_{t}^{x,\beta}}{\epsilon}$ (4.21)
be the variation process. Then it can be shown that $R$ satisfies
$\begin{cases}\mathrm{d}R_{t}&=\left(\partial_{x}b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot R_{t}+\partial_{a}b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot\gamma_{t}\right)\mathrm{d}t,\\ R_{0}&=0.\end{cases}$ (4.22)
The function $\beta\to J(\beta)$ is Gâteaux differentiable in the direction $\beta$ and its derivative is given by
[…]
Define the generalized Hamiltonian
$\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy,$ (4.24)
and introduce the adjoint process which is determined by an infinite-time BSDE
$\mathrm{d}Y_{t}^{x,\beta}=-\left(\partial_{x}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})\right)\mathrm{d}t+Z_{t}^{x,\beta}\mathrm{d}B_{t}.$ (4.25)
Applying Itô’s formula to the process $(e^{-rt}R_{t}Y_{t}^{x,\beta})$, by simple computation we can deduce that
$\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}J(\beta+\epsilon\gamma)\right|_{\epsilon=0}=\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\partial_{a}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})\cdot\gamma_{t}\mathrm{d}t\right].$ (4.26)
Thus when $\beta$ is an optimal admissible control with the associated stochastic processes $(X_{t}^{x,\beta},Y_{t}^{x,\beta},Z_{t}^{x,\beta})$, it holds that
$\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})=\min_{a\in A}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},a,Y_{t}^{x,\beta}).$ (4.27)
Recalling our convexity assumptions on $b,f$ in Assumption (4.1), we know that the representative player’s optimal control $\beta$ takes a feedback form, that is
$\beta_{t}=\hat{\alpha}(X_{t}^{x,\beta},Y_{t}^{x,\beta}).$ (4.28)
Now we consider the following two McKean–Vlasov FBSDEs:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\xi+b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.29)
$\begin{cases}\mathrm{d}X_{t}^{x}=x+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.30)
Here (4.29) and (4.30) denote the state processes of social equilibrium and the representative player, respectively. Observe that their admissible controls both take the identical feedback form $\hat{\alpha}(x,y)$. We shall prove that when social equilibrium employ this feedback control, the representative player’s loss function is minimized if they use the same feedback form—thereby constituting a Nash equilibrium.
Theorem 4.2
Let $(b,f)$ be differentiable in $(x,a)$, $\mathcal{H}$ be convex in $(x,a)$. Suppose that $\hat{\alpha}$ is Lipschitz continuous and that both (4.29) and (4.30) admit unique strong solutions in $L^{2}_{r}$. If we denote $\hat{\alpha}(X_{t}^{x},Y_{t}^{x})$ as $\alpha^{*}_{t}$ which is an admissible control in $\mathcal{A}$, then we have
$J({\alpha}^{*})=\min_{\beta\in\mathcal{A}}J(\beta).$ (4.31)
Proof. For an arbitrary admissible control $\beta\in\mathcal{A}$ and its associated process
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi}},\beta_{s})\mathrm{d}s+B_{t},$ (4.32)
we have
$J(\alpha^{*})-J(\beta)=$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})-\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x})\right)\mathrm{d}t\right]$ (4.33)
$-\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t})-b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\right)\cdot Y_{t}^{x}\mathrm{d}t\right]$
$+r\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}(X_{t}^{x}-X_{t}^{x,\beta})\cdot Y_{t}^{x}\mathrm{d}t\right].$
Since $X_{t}^{x},X_{t}^{x,\beta},Y_{t}^{x}$ are all belong to $L^{2}_{r}$, we can find a sequence of $T_{i}\rightarrow\infty$ such that
$\mathbb{E}\left[e^{-rT_{i}}(X_{T_{i}}^{x}-X_{T_{i}}^{x,\beta})\cdot Y_{T_{i}}^{x}\right]\to 0.$ (4.34)
Applying Itô’s formula to $e^{-rT_{i}}(X_{T_{i}}^{x}-X_{T_{i}}^{x,\beta})\cdot Y_{T_{i}}^{x}$ and letting $T_{i}\to\infty$, we obtain that
$\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}(X_{t}^{x}-X_{t}^{x,\beta})\left(\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})\right)\mathrm{d}t\right]$ (4.35)
$=$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(-r(X_{t}^{x}-X_{t}^{x,\beta})+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t})-b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\right)\cdot Y_{t}^{x}\mathrm{d}t\right].$
According to the convexity and differentiability of $\mathcal{H}$, we have
$\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x})-\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})$ (4.36)
$\geq$ $(X_{t}^{x,\beta}-X_{t}^{x})\cdot\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})+(\beta_{t}-{\alpha}^{*}_{t})\cdot\partial_{a}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})$
$=$ $(X_{t}^{x,\beta}-X_{t}^{x})\cdot\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x}).$
The last equality follows from the fact that $\alpha^{*}_{t}=\hat{a}(X_{t}^{x},Y_{t}^{x})$ is the minimizer of $\min_{a\in A}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},a,Y_{t}^{x})$. Combining equations (4.33), (4.35), and (4.36), we obtain
$J(\alpha^{*})-J(\beta)\leq 0$ (4.37)
for all admissible control $\beta$. Thus we complete the proof.
4.2 Solvability of mean field game FBSDEs
In this subsection, we will find sufficient conditions for the existence and uniqueness of solutions to (4.29) and (4.30). Considering the linear case, we assume $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
We require a technical lemma about the Lipschitz and convex property of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in ([1], Lemma 3.1).
Lemma 4.3
Suppose $f_{1}$ is once continuously differentiable in $(x,a)$, $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$. And $f_{1}$ is $\eta$-convex in $a$, which means
$f_{1}(x,a^{\prime})-f_{1}(x,a)-(a^{\prime}-a)\cdot\partial_{a}f_{1}(x,a)\geq\eta|a^{\prime}-a|^{2},\quad~{}\mbox{for all $x\in\mathbb{R}$.}$ (4.38)
Then it holds that
$|\hat{\alpha}(x,y)-\hat{\alpha}(x^{\prime},y^{\prime})|\leq\frac{\ell}{2\eta}|x^{\prime}-x|+\frac{|b_{3}|}{2\eta}|y^{\prime}-y|,$ (4.39)
and for some $a_{0}\in A$
$|\hat{\alpha}(x,y)|\leq\eta^{-1}(|\partial_{a}f_{1}(x,a_{0})|+|b_{2}y|)+|a_{0}|.$ (4.40)
Furthermore, if $A=\mathbb{R}$ and $\partial_{a}f$ is $\zeta$-Lipchitz in $a$, it follows that
$b_{3}(y^{\prime}-y)\cdot\left(\hat{\alpha}(x,y^{\prime})-\hat{\alpha}(x,y)\right)\leq-\frac{2b_{3}^{2}\eta}{\zeta^{2}}(y^{\prime}-y)^{2}.$ (4.41)
Theorem 4.4
Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+f_{1}(x,a)$. Under the following conditions, Assumption 3.1 and 4.1 are satisfied, and both (4.29) and (4.30) admit unique strong solutions in $L^{2}_{r}$.
(i) There exists a positive constant $k$ such that $|b_{2}|\leq k$ and $-b_{1}\geq k-\frac{r}{2}$. $f_{1}(x,a)$ is once continuously differentiable and of at most quadratic growth in $(x,a)$.
(ii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds
$f_{1}(x^{\prime},a^{\prime})-f_{1}(x,a)-\partial_{(x,a)}f_{1}(x,a)\cdot(x^{\prime}-x,a^{\prime}-a)$ (4.42)
$\geq$ $\iota(x^{\prime}-x)^{2}+\eta(a^{\prime}-a)^{2}.$
(iii) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipchitz continuous in $(x,a)$.
(iv) $A=\mathbb{R}$, and it holds that
$\min\left\{2\iota-\frac{\ell^{2}}{2\eta}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}-|b_{4}|,\frac{2b_{3}^{2}\eta}{\zeta^{2}}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.43)
Remark 4.5
If we set $f_{1}(x,a)=Ax^{2}+Ca^{2}$, we have $\iota=A,\eta=C,\zeta=2C$. Then the requirement in (4.43) becomes
$\min\left\{2A-\frac{l^{2}}{2C}-\frac{b_{3}l}{2C}-\frac{|b_{2}|}{2}-|b_{4}|,\frac{b_{3}}{2C}(b_{3}-l)-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.44)
Fixing $C,b_{2},b_{4},l,$ we take a large $b_{3}$ such that $\frac{2b_{3}^{2}\eta}{\zeta^{2}}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}$ is greater than $r/2$. Then we choose a sufficiently large $A$ such that $2A-\frac{l^{2}}{2C}-\frac{b_{3}l}{2C}-\frac{|b_{2}|}{2}-|b_{4}|$ exceeds $r/2$. This construction satisfies all the required conditions.
Proof. It’s clear that Assumption 3.1 is satisfied. And by Lemma 4.3, $\hat{\alpha}$ is Lipchitz. In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+b_{4}x\bar{\mu}+f_{1}(x,a)$ is convex in $(x,a)$ according to assumption on $f_{1}$, thus Assumption 4.1 is satisfied. To prove that the BSDEs (4.29) and (4.30) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
We consider the FBSDE (4.29), and set
$B(t,x,y,\mu)=b_{1}x+b_{2}\overline{\mu}+b_{3}\hat{\alpha}(x,y),$ (4.45)
$F(t,x,y,\mu)=b_{1}y+b_{4}\overline{\mu}+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
Take four arbitrary square integrable random variables $X,Y,X^{\prime},Y^{\prime}$. Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ We have
$-r\hat{X}\hat{Y}-\hat{X}[F(t,U)-F(t,U^{\prime})]+\hat{Y}[B(t,U)-B(t,U^{\prime})]$ (4.46)
$=$ $-r\hat{X}\hat{Y}-\hat{X}\left((b_{1}-r)\hat{Y}+\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))+b_{4}\mathbb{E}[\hat{X}]\right)$
$+\hat{Y}\left(b_{1}\hat{X}+{b}_{2}\mathbb{E}[\hat{X}]+b_{3}(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
$=$ $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
$+b_{2}\hat{Y}\mathbb{E}[\hat{X}]-b_{4}\hat{X}\mathbb{E}[\hat{X}].$
Since $f$ is $\iota$-convex in $x$, we have
$\left[\partial_{x}f_{1}(x^{\prime},a)-\partial_{x}f_{1}(x,a)\right](x^{\prime}-x)\geq 2\iota(x^{\prime}-x)^{2}.$ (4.47)
Moreover, $\partial_{x}f$ is $l$-Lipschitz in $a$ and $\hat{\alpha}$ satisfies (4.39), we have that
$-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))\right)$ (4.48)
$=$ $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))\right)$
$-\hat{X}\left(\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
$\leq$ $-2\iota\hat{X}^{2}+l|\hat{X}|(\frac{l}{2\eta}|\hat{X}|+\frac{|b_{3}|}{2\eta}|\hat{Y}|).$
From Lemma 4.3, $\hat{\alpha}$ satisfies (4.41), it follows that
$b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$ (4.49)
$=$ $b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X,Y^{\prime})\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y^{\prime})-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
$\leq-\frac{2b_{3}^{2}\eta}{\zeta^{2}}\hat{Y}^{2}+|b_{3}\hat{Y}|\left(\frac{l}{2\eta}|\hat{X}|\right).$
By applying elementary estimates, we derive
$\mathbb{E}\left[-r\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (4.50)
$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2}+|b_{4}|)\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2})\mathbb{E}[\hat{Y}^{2}]$
$<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
Now we know (4.29) admits a unique solution $X_{t}^{\xi}\in L^{2}_{r}$ and set $\bar{\mu}_{t}=\mathbb{E}[X_{t}^{\xi}]$ For the FBSDE(4.30), we set
$B^{\prime}(t,x,y)=b_{1}x+b_{2}\overline{\mu}_{t}+b_{3}\hat{\alpha}(x,y),$ (4.51)
$F^{\prime}(t,x,y)=b_{1}y+b_{4}\overline{\mu}_{t}+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
Following the identical analytical procedure, we have
$\mathbb{E}\left[-r\hat{X}\hat{Y}-\hat{X}(F^{\prime}(t,X,Y)-F^{\prime}(t,X^{\prime},Y^{\prime}))+\hat{Y}(B^{\prime}(t,X,Y)-B^{\prime}(t,X^{\prime},Y^{\prime}))\right]$ (4.52)
$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta})\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta})\mathbb{E}[\hat{Y}^{2}]$
$<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
5 Master equation representation
While we have derived a Nash equilibrium solution through FBSDEs that yields identical feedback forms for both the representative player and social equilibrium, this feedback structure differs from our previously defined formulation in Definition 3.2. In this section, we shall establish an alternative representation of the Nash equilibrium using classical solutions to the elliptic master equation (1.7).
Following Assumption 4.1, we define
$H(x,\mu,y)=H_{0}(x,\mu,\hat{\alpha}(x,y),y).$ (5.53)
Through the assumptions on $b,f$, we can easily deduce the relationship
$\partial_{y}H(x,\mu,y)=b(x,\mu,\hat{\alpha}(x,y)).$ (5.54)
Assume the master equation (1.7) has a classical solution $U(x,\mu)\in{\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ with $H$ defined above, and $\hat{b}(x,\mu)=b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$, suppose the following SDE admits a unique solution in $L^{2}_{r}$:
$\mathcal{X}_{t}^{\xi}=\xi+\int_{0}^{t}b\left(\mathcal{X}_{s}^{\xi},\mathcal{L}_{\mathcal{X}_{s}^{\xi}},\hat{\alpha}\left(\mathcal{X}_{s}^{\xi},\partial_{x}U\left(\mathcal{X}_{s}^{\xi},\mathcal{L}_{\mathcal{X}_{s}^{\xi}}\right)\right)\right)\mathrm{d}s+B_{t}.$ (5.55)
Then we take $\mathcal{Y}_{t}^{\xi}=\partial_{x}U(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}})$. By differentiating both sides of the master equation (1.7) with respect to $x$ and applying Itô’s formula, we obtain
$\begin{split}\mathrm{d}\mathcal{Y}_{t}^{\xi}=&\Bigg{[}\partial_{xx}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\cdot b\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\hat{\alpha}\left(\mathcal{X}_{t}^{\xi},\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\right)\right)\\ &+\frac{1}{2}\partial_{xx}\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)+\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\Bigg{[}\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\tilde{\mathcal{X}}^{\xi}_{t}\right)\\ &+\partial_{y}H\left(\tilde{\mathcal{X}}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\partial_{x}U\left(\tilde{\mathcal{X}}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\right)\cdot\partial_{\mu}\partial_{x}U\left(\mathcal{X}^{\xi}_{t},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\tilde{\mathcal{X}}^{\xi}_{t}\right)\Bigg{]}\Bigg{]}\mathrm{d}t\\ &+\partial_{xx}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\mathrm{d}B_{t}\\ =&\left(r\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)-\partial_{x}H\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\right)\right)\mathrm{d}t+\partial_{xx}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\mathrm{d}B_{t}.\end{split}$ (5.56)
By comparing it with (4.29), we derive the following relationship for social equilibrium:
$Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (5.57)
Applying the same argument to (4.30), we obtain:
$Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (5.58)
This demonstrates that both the representative player and social equilibrium employ the same closed-loop control
$\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu)).$ (5.59)
We now revisit the mean field games through the master equation. Let $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ be the initial state with distribution $\mu$. We prove that under the assumption that the master equation admits a classical solution, the feedback control defined by (5.59) constitutes a Nash equilibrium. Moreover, the solution to the master equation is precisely the value function of the representative player.
Theorem 5.1
Assume the master equation (1.7) admits a bounded classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ with bounded derivatives, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
$\begin{cases}X_{t}^{\xi}=\xi+\int_{0}^{t}b(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}},\hat{\alpha}(X_{s}^{\xi},\partial_{x}U(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}})))\mathrm{d}t+B_{t},\\ X_{t}^{x}=x+\int_{0}^{t}b(X_{s}^{x},\mathcal{L}_{X_{s}^{x}},\hat{\alpha}(X_{s}^{x},\partial_{x}U(X_{s}^{x},\mathcal{L}_{X_{s}^{x}})))\mathrm{d}t+B_{t}.\end{cases}$ (5.60)
Then for every initial distribution $\mu_{0}$, $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium, and $U(x,\mu_{0})$ is exactly the value function given $\alpha^{*}$.
Proof. Given $\alpha^{*}$ and initial $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu_{0})$, the states of social equilibrium is governed by the following SDE:
$X_{t}^{\xi}=\xi+\int_{0}^{t}\partial_{y}H(X_{s}^{\xi},\rho_{s},\partial_{x}U(X_{s}^{\xi},\rho_{s}))\mathrm{d}s+B_{t},\quad\quad\rho_{s}\triangleq\mathcal{L}_{X_{s}^{\xi}}.$ (5.61)
We note that this SDE is the same as (5.55). Meanwhile, the state of the representative player is governed by
$X_{t}^{x}=x+\int_{0}^{t}b(X_{s}^{x},\rho_{s},\beta_{s})\mathrm{d}s+B_{t},$ (5.62)
where $\beta\in\mathcal{A}$ remains to be determined. Applying Itô’s formula to $\mathrm{e}^{-rt}U(X^{x}_{t},\rho_{t})$ yields
$\begin{split}\mathbb{E}\mathrm{e}^{-rT}U(X^{x}_{T},\rho_{T})=&U(x,\mu_{0})+\mathbb{E}\int_{0}^{T}\big{[}-r\mathrm{e}^{-rt}U(X^{x}_{t},\rho_{t})+\mathrm{e}^{-rt}\partial_{x}U(X^{x}_{t},\rho_{t})\cdot b(X_{t}^{x},\rho_{t},\beta_{t})\\ &+\frac{1}{2}\mathrm{e}^{-rt}\partial_{xx}U(X^{x}_{t},\rho_{t})+\mathrm{e}^{-rt}\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\big{[}\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(X^{x}_{t},\rho_{t},\tilde{X}^{\xi}_{t})\\ &+\partial_{y}H(\tilde{X}_{t}^{\xi},\rho_{t},\partial_{x}U(\tilde{X}_{t}^{\xi},\rho_{t}))\cdot\partial_{\mu}U(X^{x}_{t},\rho_{t},\tilde{X}^{\xi}_{t})\big{]}\big{]}\mathrm{d}t\\ =&U(x,\mu_{0})+\mathbb{E}\int_{0}^{T}\mathrm{e}^{-rt}\left[\partial_{x}U(X^{x}_{t},\rho_{t})\cdot b(X_{t}^{x},\rho_{t},\beta_{t})-H(X^{x}_{t},\rho_{t},\partial_{x}U(X^{x}_{t},\rho_{t}))\right]\mathrm{d}t\\ \geq&U(x,\mu_{0})-\mathbb{E}\int_{0}^{T}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\beta_{t})\mathrm{d}t.\end{split}$ (5.63)
Since $U$ and its derivatives are bounded, taking the limit as $T\rightarrow+\infty$ yields
$U(x,\mu_{0})\leq\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\beta_{t})\mathrm{d}t$ (5.64)
for every feasible control $\beta$, and the equality holds when $\beta_{t}=\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t}))$. This shows $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is an Nash equilibrium. And we have
$U(x,\mu_{0})=\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t})))\mathrm{d}t,$ (5.65)
this completes the proofs of our desired results.

### v2

@@
 We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations.
-By employing the elliptic master equation, we provide a representation for the Nash equilibrium.
-Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, elliptic master equation.
+By employing the elliptic master equations, a class of distribution-dependent elliptic PDEs , we provide a representation for the Nash equilibrium.
+And we prove that the solutions to a system of infinite-time forward-backward stochastic differential equations can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
+Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, elliptic master equations.
 1 Introduction
@@
 The dynamic of the private state of the representative player is given by
-$X^{x}_{t}=x+\int_{0}^{t}b(X^{x}_{t},\mu_{s},\beta_{s}^{x})\mathrm{d}t+B_{t}.$ (1.1)
+$X^{x}_{t}=x+\int_{0}^{t}b(X^{x}_{t},\mu_{s},\beta_{s}^{x})\mathrm{d}s+B_{t}.$ (1.1)
 Here $\mu_{t}$ is the population distribution , while $\beta^{x}$ is the optimal strategy of the representative player, derived as the solution to the stochastic optimization problem:
@@
 We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time forward-backward stochastic differential equations.
-The foundational work in [14] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, later extended by [1] to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games.
+The foundational work in [14] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, and subsequent work in [15] investigated a more general setting and established connections with quasilinear elliptic PDEs.
+Recently, [1] extended this framework to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games.
 In this paper, we partition all agents into a representative player and the social equilibrium, and characterize the equilibrium state through the following system of infinite-time FBSDE:
-$\begin{cases}\mathrm{d}X_{t}^{\xi}=\xi+b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.5)
-$\begin{cases}\mathrm{d}X_{t}^{x}=x+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.6)
+$\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.5)
+$\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.6)
 Here $\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy$ is the generalized Hamiltonian, and $\hat{\alpha}(x,y)$ is the minimizer of $\mathcal{H}(x,\mu,a,y)$ with respect to $a$ when we assume $\mathcal{H}$ is separable in variables $\mu$ and $a$.
@@
 This is precisely the relationship expressed in Equation (1.4).
+This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
 2 Preliminaries
@@
 For any sub-$\sigma$-field ${\cal G}\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, ${\cal G}$-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ For any $\mu,\nu\in{\cal P}_{p}$, we define the $W_{p}$–Wasserstein distance between them as follows:
-$W_{p}(\mu,\nu)\triangleq\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~{}\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
+$W_{p}(\mu,\nu):=\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~{}\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
 Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}-$valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
@@
 $J\left(x,\xi;\alpha,\beta\right)=\mathbb{E}\bigg{[}\int_{0}^{+\infty}\mathrm{e}^{-rt}f\big{(}X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi,\alpha}},\beta_{t}\big{)}\mathrm{d}t\bigg{]}.$ (3.8)
-We observe that the cost function depends exclusively on the distribution of $X_{t}^{\xi,\alpha}$ and we have
-$\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad J(x,\xi^{\prime};\alpha,\beta)=J(x,\xi;\alpha,\beta),\quad~{}\mbox{for all $x,\alpha,\beta$.}$ (3.9)
-Therefore, we may define
-$J(x,\mu;\alpha,\beta)\triangleq J(x,\xi;\alpha,\beta),\quad\quad\xi\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.10)
-Now, for any $(x,\mu)\in\mathbb{R}\times{\cal P}_{2}$ and $\alpha\in\mathcal{A}$, we consider the infimum
-$V(x,\mu;\alpha)\triangleq\inf_{\beta\in\mathcal{A}}J(x,\mu;\alpha,\beta).$ (3.11)
+For any $(x,\xi)\in\mathbb{R}\times\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$, we consider the infimum
+$V(x,\xi;\alpha)\triangleq\inf_{\beta\in\mathcal{A}}J(x,\xi;\alpha,\beta).$ (3.9)
 Definition 3.2
-We say a function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if the closed-loop controls $\alpha_{s}^{\mu_{0}}={\alpha}^{*}(X_{s}^{\xi,\alpha^{\mu_{0}}},\mathcal{L}_{X_{s}^{\xi,\alpha^{\mu_{0}}}})$,$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi,\alpha^{\mu_{0}}}})$ satisfy
-$\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\mu_{0};\alpha^{\mu_{0}},\beta).$ (3.12)
-When this Nash equilibrium $\alpha^{*}$ exists, we can define the value function with respect to this Nash equilibrium as:
-$V(x,\mu_{0})\triangleq J(x,\mu_{0};\alpha^{\mu_{0}},\alpha^{x}).$ (3.13)
+We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$,$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
+$\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\xi_{0};\alpha^{\xi_{0}},\beta).$ (3.10)
+When this Nash equilibrium $\alpha^{*}$ exists, we have
+$\xi,\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad\mathcal{L}_{X_{t}^{\xi,\alpha^{\xi}}}=\mathcal{L}_{X_{t}^{\xi^{\prime},\alpha^{\xi^{\prime}}}},\quad~{}\mbox{for a.e.
+$t\geq 0$.}$ (3.11)
+Therefore, we can define
+$V(x,\mu)\triangleq J(x,\xi_{0};\alpha^{\xi_{0}},\alpha^{x}),\quad\quad\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.12)
 Note that our definition of Nash equilibrium differs from that in [1] and earlier finite-time mean field games.
@@
 Now we consider the following two McKean–Vlasov FBSDEs:
-$\begin{cases}\mathrm{d}X_{t}^{\xi}=\xi+b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.29)
-$\begin{cases}\mathrm{d}X_{t}^{x}=x+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.30)
+$\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.12)
+$\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.13)
 Here (4.12) and (4.13) denote the state processes of social equilibrium and the representative player, respectively.
@@
 Remark 4.5
-If we set $f_{1}(x,a)=Ax^{2}+Ca^{2}$, we have $\iota=A,\eta=C,\zeta=2C$. Then the requirement in (4.43) becomes
+If we set $f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C$. Then the requirement in (4.26) becomes
 $\min\left\{2A-\frac{l^{2}}{2C}-\frac{b_{3}l}{2C}-\frac{|b_{2}|}{2}-|b_{4}|,\frac{b_{3}}{2C}(b_{3}-l)-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.27)
@@
 Theorem 5.1
-Assume the master equation (1.7) admits a bounded classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ with bounded derivatives, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
+Assume the master equation (1.7) admits a classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
 $\begin{cases}X_{t}^{\xi}=\xi+\int_{0}^{t}b(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}},\hat{\alpha}(X_{s}^{\xi},\partial_{x}U(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}})))\mathrm{d}t+B_{t},\\ X_{t}^{x}=x+\int_{0}^{t}b(X_{s}^{x},\mathcal{L}_{X_{s}^{x}},\hat{\alpha}(X_{s}^{x},\partial_{x}U(X_{s}^{x},\mathcal{L}_{X_{s}^{x}})))\mathrm{d}t+B_{t}.\end{cases}$ (5.8)
@@
 $\begin{split}\mathbb{E}\mathrm{e}^{-rT}U(X^{x}_{T},\rho_{T})=&U(x,\mu_{0})+\mathbb{E}\int_{0}^{T}\big{[}-r\mathrm{e}^{-rt}U(X^{x}_{t},\rho_{t})+\mathrm{e}^{-rt}\partial_{x}U(X^{x}_{t},\rho_{t})\cdot b(X_{t}^{x},\rho_{t},\beta_{t})\\ &+\frac{1}{2}\mathrm{e}^{-rt}\partial_{xx}U(X^{x}_{t},\rho_{t})+\mathrm{e}^{-rt}\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\big{[}\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(X^{x}_{t},\rho_{t},\tilde{X}^{\xi}_{t})\\ &+\partial_{y}H(\tilde{X}_{t}^{\xi},\rho_{t},\partial_{x}U(\tilde{X}_{t}^{\xi},\rho_{t}))\cdot\partial_{\mu}U(X^{x}_{t},\rho_{t},\tilde{X}^{\xi}_{t})\big{]}\big{]}\mathrm{d}t\\ =&U(x,\mu_{0})+\mathbb{E}\int_{0}^{T}\mathrm{e}^{-rt}\left[\partial_{x}U(X^{x}_{t},\rho_{t})\cdot b(X_{t}^{x},\rho_{t},\beta_{t})-H(X^{x}_{t},\rho_{t},\partial_{x}U(X^{x}_{t},\rho_{t}))\right]\mathrm{d}t\\ \geq&U(x,\mu_{0})-\mathbb{E}\int_{0}^{T}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\beta_{t})\mathrm{d}t.\end{split}$ (5.11)
-Since $U$ and its derivatives are bounded, taking the limit as $T\rightarrow+\infty$ yields
+Since $U$ is of at most quadratic growth, taking the limit as $T\rightarrow+\infty$ yields
 $U(x,\mu_{0})\leq\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\beta_{t})\mathrm{d}t$ (5.12)
@@
 this completes the proofs of our desired results.
+At the end of this section, we provide an example of a solvable elliptic master equation.
+Set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a,f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, where $b_{1},b_{2},b_{3},b_{4},A,C$ are constants satisfying all conditions in Theorem 4.4 and Remark 4.5.
+Then we have $\hat{\alpha}(x,y)=-\frac{b_{3}y}{2C}$ and
+$H(x,\mu,y)=(b_{1}x+b_{2}\bar{\mu})\cdot y+b_{4}x\bar{\mu}+Ax^{2}-\frac{b_{3}^{2}}{4C}y^{2}.$ (5.14)
+The master equation (1.7) now becomes
+$\begin{split}rU(x,\mu)=&(b_{1}x+b_{2}\bar{\mu})\cdot\partial_{x}U(x,\mu)+b_{4}x\bar{\mu}+Ax^{2}\\ &-\frac{b_{3}^{2}}{4C}(\partial_{x}U(x,\mu))^{2}+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})(b_{1}\tilde{\xi}+b_{2}\bar{\mu}-\frac{b^{2}_{3}}{2C}\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (5.15)
+Theorem 5.2
+The master equation (5.15) admits a classical solution $U(x,\mu)$ in ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$.
+Proof. We assume that the solution takes the form
+$U(x,\mu)=a_{1}x^{2}+a_{2}x\bar{\mu}+a_{3}(\bar{\mu})^{2}+a_{4},$ (5.16)
+where $a_{1},a_{2},a_{3},a_{4}$ are constants remain to be determined. Then we have
+$\begin{split}&\partial_{x}U(x,\mu)=2a_{1}x+a_{2}\bar{\mu},\quad\quad\quad\partial_{xx}U(x,\mu)=2a_{1},\\ &\partial_{\mu}U(x,\mu,\tilde{x})=a_{2}x+2a_{3}\bar{\mu},\quad\quad\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{x})=0.\end{split}$ (5.17)
+Substituting these into the equation (5.15), we obtain
+$\begin{split}&r\left(a_{1}x^{2}+a_{2}x\bar{\mu}+a_{3}(\bar{\mu})^{2}+a_{4}\right)\\ =&(b_{1}x+b_{2}\bar{\mu})\cdot(2a_{1}x+a_{2}\bar{\mu})+b_{4}x\bar{\mu}+Ax^{2}-\frac{b_{3}^{2}}{4C}(2a_{1}x+a_{2}\bar{\mu})^{2}+a_{1}\\ &+(a_{2}x+2a_{3}\bar{\mu})\cdot\left(b_{1}\bar{\mu}+b_{2}\bar{\mu}-\frac{b^{2}_{3}}{2C}(2a_{1}\bar{\mu}+a_{2}\bar{\mu})\right).\end{split}$ (5.18)
+Comparing all coefficients, we get the following system of linear and quadratic equations:
+$\begin{cases}ra_{1}=2b_{1}a_{1}+A-\frac{b_{3}^{2}}{C}a_{1}^{2},\\ ra_{2}=2b_{2}a_{1}+b_{1}a_{2}+b_{4}-\frac{b_{3}^{2}}{C}a_{1}a_{2}+a_{2}(b_{1}+b_{2}-\frac{b_{3}^{2}}{C}a_{1}-\frac{b_{3}^{2}}{2C}a_{2}),\\ ra_{3}=b_{2}a_{2}-\frac{b_{3}^{2}}{4C}a_{2}^{2}+2a_{3}(b_{1}+b_{2}-\frac{b_{3}^{2}}{C}a_{1}-\frac{b_{3}^{2}}{2C}a_{2}),\\ ra_{4}=a_{1}.\end{cases}$ (5.19)
+Then, we can solve for $a_{1},a_{2},a_{3},a_{4}$.
+Remark 5.3
+We have
+$\hat{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))=(b_{1}-\frac{a_{1}b_{3}^{2}}{C})x+(b_{2}-\frac{a_{2}b_{3}^{2}}{2C})\bar{\mu}.$ (5.20)
+Since $b_{1}+|b_{2}|<\frac{r}{2}$, if we further require $a_{1}>0,a_{2}>0$, the SDE (5.3) admits a unique solution in $L_{r}^{2}$.
+6 Viscosity solution to distribution-dependent elliptic PDE
+6.1 Regularity of the FBSDE solutions
+We have proven that the classical solutions of the master equations can be employed to resolve the infinite-time FBSDEs.
+In this section, the solution of the infinite-time FBSDEs will be utilized to characterize the viscosity solutions of the distribution-dependent elliptic PDEs.
+Specifically, we consider the following FBSDEs with initial state $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $x\in\mathbb{R}$:
+$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.1)
+$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (6.2)
+Here $H(x,\mu,y)$ is defined in (5.1) and the above equations are the same as (1.5) and (1.6).
+We define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ and attempt to prove that it satisfy the following elliptic PDE:
+$\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (6.3)
+This equation is derived by taking the partial derivative of both sides of the master equation (1.7) with respect to $x$. We have the relationship $\mathcal{U}(x,\mu)=\partial_{x}U(x,\mu)$.
+The proof of $\mathcal{V}$ being a viscosity solution to equation (6.3) rests on two fundamental results:
+  • The value of $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$, but not on the specific realization of $\xi$.
+  • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on ×$\mathbb{R}\times\mathcal{P}_{2}$.
+While the validity of these two conditions has not been established for general cases, we can prove them for linear cases.
+If we set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, and we further assume that all assumptions in Theorem 4.4 and (4.27) are satisfied, the FBSDEs (6.1) and (6.2) become the following linear equations:
+$\begin{cases}\mathrm{d}X_{t}^{\xi}=\left(b_{1}X_{t}^{\xi}-\frac{b^{2}_{3}}{2C}Y_{t}^{\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[(b_{1}-r)Y_{t}^{\xi}+2AX_{t}^{\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.4)
+$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\left(b_{1}X_{t}^{x,\xi}-\frac{b^{2}_{3}}{2C}Y_{t}^{x,\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[(b_{1}-r)Y_{t}^{x,\xi}+2AX_{t}^{x,\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right]\mathrm{d}t+Z_{t}^{x,\xi}\mathrm{d}B_{t},\\ X_{0}^{x,\xi}=x.\end{cases}$ (6.5)
+We have already proved that for any initial state $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $x\in\mathbb{R}$, equations (6.4) and (6.5) have unique solutions in $L^{2}_{r}$.
+Lemma 6.1
+For two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution, their corresponding solutions to (6.4) have the same expectations. That is
+$\mathbb{E}[X_{t}^{\xi_{1}}]=\mathbb{E}[X_{t}^{\xi_{2}}],\quad~{}\mbox{for a.e.$t\geq 0.$}$ (6.6)
+Proof.
+Fix $\xi_{1}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and the corresponding solution $X^{\xi_{1}}_{t}$ to equation (6.4).
+Taking $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]$, we consider the following two infinite-time FBSDEs:
+$\begin{cases}\mathrm{d}X_{t}^{\xi}=\left(pX_{t}^{\xi}-qY_{t}^{\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left(uY_{t}^{\xi}+vX_{t}^{\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.7)
+and
+$\begin{cases}\mathrm{d}\tilde{X}_{t}^{\xi}=\left(p\tilde{X}_{t}^{\xi}-q\tilde{Y}_{t}^{\xi}+b_{2}\phi_{t}\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}\tilde{Y}_{t}^{\xi}=-\left(u\tilde{Y}_{t}^{\xi}+v\tilde{X}_{t}^{\xi}+b_{4}\phi_{t}\right)\mathrm{d}t+\tilde{Z}_{t}^{\xi}\mathrm{d}B_{t},\\ \tilde{X}_{0}^{\xi}=\xi.\end{cases}$ (6.8)
+Where $p=b_{1},q=\frac{b^{2}_{3}}{2C},u=b_{1}-r,v=2A$, and according to the argument in Theorem 4.4, each of the aforementioned equations admits a unique solution in $L^{2}_{r}$.
+We denote the solutions for Equation (6.7) with initial states ($\xi_{1},\xi_{2}$) by ($X_{t}^{\xi_{1}},X_{t}^{\xi_{2}}$), and that for Equation (6.8) under identical initial states by ($\tilde{X}_{t}^{\xi_{1}},\tilde{X}_{t}^{\xi_{2}}$).
+For Equation (6.8), let $P_{t}^{\xi}=m\tilde{X}_{t}^{\xi}-\tilde{Y}_{t}^{\xi}$ where $m$ is the positive solution of $qm^{2}-(p+u)m-v=0$. Then
+$\begin{split}\mathrm{d}P_{t}^{\xi}&=\left[(mp+v)\tilde{X}_{t}^{\xi}-(mq-u)\tilde{Y}_{t}^{\xi}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+\mathrm{d}B_{t}\\ &=\left[(mq-u)P_{t}^{\xi}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+(m-\tilde{Z}_{t}^{\xi})\mathrm{d}B_{t}.\end{split}$ (6.9)
+This demonstrates that $P^{\xi}$ satisfies an infinite-time backward stochastic differential equation
+$\mathrm{d}P_{t}=\left[(mq-u)P_{t}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+Z_{t}\mathrm{d}B_{t}$ (6.10)
+and exhibits no dependence on $\xi$.
+Applying ([14], Theorem 4) with the facts that $mq-u\geq-u=r-b_{1}\geq\frac{r}{2}$ and $\phi_{t}\in L^{2}_{r}$, the above equation admits a unique solution $(P,Z)\in L^{2}_{r}$.
+Then we know $\tilde{X}_{t}^{\xi}$ satisfy the following SDE,
+$\mathrm{d}\tilde{X}_{t}^{\xi}=\left[(p-mq)\tilde{X}_{t}^{\xi}+qP_{t}+b_{2}\phi_{t}\right]\mathrm{d}t+\mathrm{d}B_{t},\quad\quad\tilde{X}_{0}^{\xi}=\xi.$ (6.11)
+It is easy to see that $\tilde{X}_{t}^{\xi_{1}}$ and $\tilde{X}_{t}^{\xi_{2}}$ have the same distribution, so we have $\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{1}}].$
+Since $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]$, $X_{t}^{\xi_{1}}$ and $\tilde{X}_{t}^{\xi_{1}}$ are the same.
+Then we have $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{1}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]$, which means $\tilde{X}_{t}^{\xi_{2}}$ is also a solution to FBSDE (6.7) with initial $\xi_{2}$.
+By the uniqueness of the solution, we get that $\mathbb{E}[{X}_{t}^{\xi_{2}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]=\mathbb{E}[X_{t}^{\xi_{1}}]$.
+Now we finish the proof.
+Remark 6.2
+For any two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution and $x\in\mathbb{R}$, the solutions $Y^{x,\xi_{1}},Y^{x,\xi_{2}}$ to (6.5) are the same.
+We can say that $Y^{x,\xi}_{0}$ depends depends solely on the distribution of $\xi$.
+Lemma 6.3
+For FBSDEs (6.4) and (6.5), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$.
+Then we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
+$|\mathcal{V}(x,\mu)-\mathcal{V}(x^{\prime},\mu^{\prime})\leq C(|x-x^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime}))$ (6.12)
+Proof.
+Let $(X^{\xi_{1}},Y^{\xi_{1}},X^{x_{1},\xi_{1}},Y^{x_{1},\xi_{1}})$ and $(X^{\xi_{2}},Y^{\xi_{2}},X^{x_{2},\xi_{2}},Y^{x_{2},\xi_{2}})$ be the solutions of Equations (6.4) and (6.5) with $\mathcal{L}_{\xi_{1}}=\mu_{1},\mathcal{L}_{\xi_{2}}=\mu_{2}$.
+Set
+$\begin{split}&\hat{X}^{\xi}=X^{\xi_{1}}-X^{\xi_{2}},\quad\hat{Y}^{\xi}=Y^{\xi_{1}}-Y^{\xi_{2}},\\ &\hat{X}^{x,\xi}=X^{x_{1},\xi_{1}}-X^{x_{2}\xi_{2}},\quad\hat{Y}^{x,\xi}=Y^{x_{1},\xi_{1}}-Y^{x_{2},\xi_{2}}.\end{split}$ (6.13)
+$C_{1},C_{2},C_{3},C_{4}$ appeared in the following proof are positive constants.
+Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{x_{1},\xi_{1}}_{t}-Y^{x_{2},\xi_{2}}_{t}|^{2}$, we get
+$\begin{split}|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}&\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\mathbb{E}\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t\\ &\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t.\end{split}$ (6.14)
+Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{x,\xi}\hat{Y}^{x,\xi}$, we get
+$\begin{split}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}\right]\mathrm{d}t\leq C_{2}\left(\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}+\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\right).\end{split}$ (6.15)
+It turns out that
+$\begin{split}(\hat{Y}^{x,\xi}_{0})^{2}\leq&C_{1}C_{2}\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}+(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\\ \leq&\frac{1}{2}(\hat{Y}^{x,\xi}_{0})^{2}+\frac{C_{1}^{2}C_{2}^{2}}{2}(\hat{X}^{x,\xi}_{0})^{2}+(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t.\end{split}$ (6.16)
+Thus
+$(\hat{Y}^{x,\xi}_{0})^{2}\leq C_{1}^{2}C_{2}^{2}(\hat{X}^{x,\xi}_{0})^{2}+2(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t.$ (6.17)
+Applying the same arguments to $\mathrm{e}^{-rt}|\hat{Y}^{\xi}_{t}|^{2}$ and $\mathrm{e}^{-rt}\hat{X}^{\xi}\hat{Y}^{\xi}$, we can get that
+$\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\leq C_{3}\mathbb{E}|\xi_{1}-\xi_{2}|^{2}.$ (6.18)
+So we can deduce that
+$|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathbb{E}|\xi_{1}-\xi_{2}|^{2}\right).$ (6.19)
+Since $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$, we have
+$|\mathcal{V}(x_{1},\mu_{1})-\mathcal{V}(x_{2},\mu_{2})|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathcal{W}_{2}^{2}(\mu_{1},\mu_{2})\right),$ (6.20)
+then we get the desired result.
+6.2 Connection with distribution-dependent elliptic PDE
+Now let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
+$(\mathcal{L}\mathcal{U})(x,\mu,\mathcal{U}(x,\mu))+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.21)
+where
+$\begin{split}(\mathcal{L}\Phi)(x,\mu,\Psi)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.22)
+and
+$F(x,\mu,\Psi)\triangleq\partial_{x}H(x,\mu,\Psi(x,\mu))-r\Psi(x,\mu).$ (6.23)
+Definition 6.4
+Let $\mathcal{U}\in C(\mathbb{R}\times\mathcal{P}_{2})$.
+Then $\mathcal{U}$ is called a viscosity subsolution (resp.
+supersolution) of PDE (6.21) if, whenever $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp.
+minimum) of $\mathcal{U}-\Psi$, we have
+$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.24)
+(respectively
+$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.25)
+). The function $\mathcal{U}$ is called a viscosity solution of PDE (6.21) if it is both a viscosity subsolution and a viscosity supersolution.
+We now assert that
+Theorem 6.5
+Assume that both FBSDEs (6.1) and (6.2) admit unique solutions in $L^{2}_{r}$, $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$.
+Then we define the function $\mathcal{V}(x,\mu)\triangleq Y_{0}^{x,\xi}$ with $\mathcal{L}_{\xi}=\mu$, and assume $\mathcal{V}$ is jointly continuous on ×$\mathbb{R}\times\mathcal{P}_{2}$.
+The function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.21).
+Proof. We only show that $\mathcal{V}$ is a viscosity subsolution of PDE (6.21). A similar argument will show that it is also a viscosity supersolution of (6.21).
+Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$.
+Let $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum of $\mathcal{U}-\Psi$.
+We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$.
+We suppose that
+$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.26)
+It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
+$\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)(x,\mu,\mathcal{V}(x,\mu))+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.27)
+Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2).
+For some $T>0$, let $\tau$ denote the stopping time
+$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\mathcal{L}_{X_{t}^{\xi^{0}}})\in O\}\land T.$ (6.28)
+We first note that the pair of processes
+$(\bar{Y}\left(t\right),\bar{Z}\left(t\right))\triangleq(Y^{x^{0},\xi^{0}}_{t\wedge\tau},I_{\left[0,\tau\right]}(t)Z^{x^{0},\xi^{0}}_{t}),\quad 0\leq t\leq T,$ (6.29)
+is the solution of BSDE
+$\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.30)
+Next, it follows from Itô’s formula that the pair of processes
+$(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}))$ (6.31)
+is the solution of BSDE
+$\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.32)
+Define
+$\beta_{s}=-\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))-F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))$ (6.33)
+and
+$\left(\tilde{Y}\left(t\right),\tilde{Z}\left(t\right)\right)=\left(\hat{Y}\left(t\right)-\bar{Y}\left(t\right),\hat{Z}\left(t\right)-\bar{Z}\left(t\right)\right).$ (6.34)
+We have
+$\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.35)
+Therefore
+$\tilde{Y}_{0}=\mathbb{E}\left[\tilde{Y}_{\tau}+\int_{0}^{\tau}\beta_{s}\mathrm{d}s\right]$ (6.36)
+Now from the choice of $O$ and $\tau$, a.s.
+$\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.37)
+Consequently, $\tilde{Y}_{0}=\Psi(x^{0},\mu^{0})-\mathcal{V}(x^{0},\mu^{0})>0$, which contradicts the earlier assumption.

### v3

@@
 When the optimal strategy of the representative player shares the same feedback form with the strategy of the social equilibrium, we say the system achieves a Nash equilibrium.
-We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations.
+We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations(FBSDEs).
 By employing the elliptic master equations, a class of distribution-dependent elliptic PDEs , we provide a representation for the Nash equilibrium.
-And we prove that the solutions to a system of infinite-time forward-backward stochastic differential equations can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
-Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, elliptic master equations.
+We prove the Yamada-Watanabe theorem and show the weak uniqueness for infinite-time FBSDEs.
+And we prove that the solutions to a system of infinite-time FBSDEs can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
+Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, weak uniqueness, elliptic master equations.
 1 Introduction
@@
 This is precisely the relationship expressed in Equation (1.4).
-This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
+This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we prove the Yamada-Watanabe theorem for infinite-time FBSDEs and provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
 2 Preliminaries
@@
 Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$, define the larger filtered probability space by
-$\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}.$ (2.1)
-Throughout the paper we will use the probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$.
-However, when we deal with the distribution-dependent master equation, independent copies of random variables or processes are needed.
-Then we will tacitly use their extensions to the larger space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}},\tilde{\mathbb{F}})$.
+$\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}$.
+(2.1)
+[…]
 Let ${\cal P}\triangleq{\cal P}(\mathbb{R})$ be the set of all probability measures on $\mathbb{R}$ and let ${\cal P}_{p}(p\geq 1)$ denote the set of $\mu\in{\cal P}$ with finite $p$-th moment.
-For any sub-$\sigma$-field ${\cal G}\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, ${\cal G}$-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ For any $\mu,\nu\in{\cal P}_{p}$, we define the $W_{p}$–Wasserstein distance between them as follows:
+For any sub-$\sigma$-field $\cal G$$\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, $\cal G$}-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ .
+For any $\mu,\nu\in{\cal P}_{p}$, we define the $W_{p}$–Wasserstein distance between them as follows:
 $W_{p}(\mu,\nu):=\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
 Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}$-valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
-[…]
+$\mathbb{E}\left[\int_{t_{0}}^{\infty}e^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
 To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$. For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
@@
 For ${\cal C}^{1}({\cal P}_{2})$, we mean the space of functions $U\in{\cal C}^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.6).
-Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on ×$\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on ×$\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
-Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on ×$\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}^{d}\times{\cal P}_{2}\times\mathbb{R}$.
+Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
+Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}^{d}\times{\cal P}_{2}\times\mathbb{R}$.
 Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5].
@@
 $\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big{[}\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\hbox{\rm tr$\,$}\big{(}\partial_{xx}U[\sigma_{t}^{1}]^{2}\big{)}\Big{]}(t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big{(}\tilde{\mathbb{E}}_{{\cal F}_{t}}\big{[}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big{]}\Big{)}\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
-Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$
+Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ .
 3 Infinite-time mean field games with discounting
@@
 Definition 3.2
-We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$,$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
+We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
 $\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\xi_{0};\alpha^{\xi_{0}},\beta).$ (3.10)
@@
 (i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$.
-Where $b_{0},f_{0}$ are measurable functions on ×$\mathbb{R}\times{\cal P}_{2}$, $b_{1},f_{1}$ are measurable functions on ×$\mathbb{R}\times A$.
+Where $b_{0},f_{0}$ are measurable functions on $\mathbb{R}\times{\cal P}_{2}$, $b_{1},f_{1}$ are measurable functions on $\mathbb{R}\times A$.
 (ii) $b,f$ are differentiable with respect to $(x,a)$ and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
@@
 Let us formally derive the maximum principle for the infinite-time control problem.
-Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ Let
+Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ .
+Let
 $R_{t}=\lim_{\epsilon\to 0}\frac{X_{t}^{x,\beta+\epsilon\gamma}-X_{t}^{x,\beta}}{\epsilon}$ (4.4)
@@
 The function $\beta\to J(\beta)$ is Gâteaux differentiable in the direction $\beta$ and its derivative is given by
-[…]
+$\left.\frac{d}{d\epsilon}J(\beta+\epsilon\gamma)\right|_{\epsilon=0}=\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(\partial_{x}f(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot R_{t}+\partial_{a}f(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot\gamma_{t}\right)dt\right].$ (4.6)
 Define the generalized Hamiltonian
@@
 In this subsection, we will find sufficient conditions for the existence and uniqueness of solutions to (4.12) and (4.13).
-Considering the linear case, we assume $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
+Considering the linear case, we assume $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
 We require a technical lemma about the Lipschitz and convex property of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in ([1], Lemma 3.1).
@@
 Theorem 4.4
-Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+f_{1}(x,a)$.
+Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$.
 Under the following conditions, Assumption 3.1 and 4.1 are satisfied, and both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
-(i) There exists a positive constant $k$ such that $|b_{2}|\leq k$ and $-b_{1}\geq k-\frac{r}{2}$. $f_{1}(x,a)$ is once continuously differentiable and of at most quadratic growth in $(x,a)$.
-(ii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds
+(i) There exists a positive constant $k$ such that $|b_{2}|\leq k$ and $b_{1}\geq k-\frac{r}{2}$.
+$f_{0}(x,\mu)$ is once continuously differentiable in $x$ and of at most quadratic growth in $(x,\mu)$.
+$f_{1}(x,a)$ is once continuously differentiable and of at most quadratic growth in $(x,a)$.
+(ii) $f_{0}(x,\mu)$ is convex in $x$ and there exists a positive constant $b_{4}>0$, such that
+$|\partial_{x}f_{0}(x,\mu)-\partial_{x}f_{0}(x^{\prime},\mu^{\prime})|<b_{4}\left(|x-x^{\prime}|+W_{2}(\mu,\mu^{\prime})\right).$ (4.25)
+(iii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds
 $f_{1}(x^{\prime},a^{\prime})-f_{1}(x,a)-\partial_{(x,a)}f_{1}(x,a)\cdot(x^{\prime}-x,a^{\prime}-a)$ (4.26)
 $\geq$ $\iota(x^{\prime}-x)^{2}+\eta(a^{\prime}-a)^{2}.$
-(iii) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipchitz continuous in $(x,a)$.
-(iv) $A=\mathbb{R}$, and it holds that
-$\min\left\{2\iota-\frac{\ell^{2}}{2\eta}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}-|b_{4}|,\frac{2b_{3}^{2}\eta}{\zeta^{2}}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.26)
+(iv) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipchitz continuous in $(x,a)$.
+(v) $A=\mathbb{R}$, and it holds that
+$\min\left\{2\iota-\frac{\ell^{2}}{2\eta}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}-2b_{4},\frac{2b_{3}^{2}\eta}{\zeta^{2}}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.27)
 Remark 4.5
-If we set $f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C$. Then the requirement in (4.26) becomes
-$\min\left\{2A-\frac{l^{2}}{2C}-\frac{b_{3}l}{2C}-\frac{|b_{2}|}{2}-|b_{4}|,\frac{b_{3}}{2C}(b_{3}-l)-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.27)
-Fixing $C,b_{2},b_{4},l,$ we take a large $b_{3}$ such that $\frac{2b_{3}^{2}\eta}{\zeta^{2}}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}$ is greater than $r/2$.
-Then we choose a sufficiently large $A$ such that $2A-\frac{l^{2}}{2C}-\frac{b_{3}l}{2C}-\frac{|b_{2}|}{2}-|b_{4}|$ exceeds $r/2$.
+If we set $f_{0}(x,\mu)=b_{4}x\bar{\mu},f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C,\ell=0$. Then the requirement in (4.27) becomes
+$\min\left\{2A-\frac{|b_{2}|}{2}-|b_{4}|,\frac{b_{3}^{2}}{2C}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.28)
+Fixing $C,b_{2},b_{4},$ we take a large $b_{3}$ such that $\frac{b_{3}^{2}}{2C}-\frac{|b_{2}|}{2}$ is greater than $r/2$.
+Then we choose a sufficiently large $A$ such that $2A-\frac{|b_{2}|}{2}-|b_{4}|$ exceeds $r/2$.
 This construction satisfies all the required conditions.
@@
 And by Lemma 4.3, $\hat{\alpha}$ is Lipchitz.
-In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+b_{4}x\bar{\mu}+f_{1}(x,a)$ is convex in $(x,a)$ according to assumption on $f_{1}$, thus Assumption 4.1 is satisfied.
+In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+f_{0}(x,\mu)+f_{1}(x,a)$ is convex in $(x,a)$ according to assumption on $f_{0},f_{1}$, thus Assumption 4.1 is satisfied.
 To prove that the BSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
@@
 $B(t,x,y,\mu)=b_{1}x+b_{2}\overline{\mu}+b_{3}\hat{\alpha}(x,y),$ (4.29)
-$F(t,x,y,\mu)=b_{1}y+b_{4}\overline{\mu}+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
+$F(t,x,y,\mu)=b_{1}y+\partial_{x}f_{0}(x,\mu)+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
 Take four arbitrary square integrable random variables $X,Y,X^{\prime},Y^{\prime}$.
-Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ We have
+Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ .
+We have
 $-r\hat{X}\hat{Y}-\hat{X}[F(t,U)-F(t,U^{\prime})]+\hat{Y}[B(t,U)-B(t,U^{\prime})]$ (4.30)
-$=$ $-r\hat{X}\hat{Y}-\hat{X}\left((b_{1}-r)\hat{Y}+\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))+b_{4}\mathbb{E}[\hat{X}]\right)$
+$=$ $-r\hat{X}\hat{Y}-\hat{X}\left((b_{1}-r)\hat{Y}+\partial_{x}f_{0}(X,\mathcal{L}_{X})-\partial_{x}f_{0}(X^{\prime},\mathcal{L}_{X^{\prime}})+\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
 $+\hat{Y}\left(b_{1}\hat{X}+{b}_{2}\mathbb{E}[\hat{X}]+b_{3}(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
-$=$ $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
-$+b_{2}\hat{Y}\mathbb{E}[\hat{X}]-b_{4}\hat{X}\mathbb{E}[\hat{X}].$
-Since $f$ is $\iota$-convex in $x$, we have
+$\leq$ $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
+$+b_{2}\hat{Y}\mathbb{E}[\hat{X}]+b_{4}|\hat{X}|\cdot\left(|\hat{X}|+(\mathbb{E}[\hat{X}^{2}])^{1/2}\right).$
+Since $f_{1}$ is $\iota$-convex in $x$, we have
 $\left[\partial_{x}f_{1}(x^{\prime},a)-\partial_{x}f_{1}(x,a)\right](x^{\prime}-x)\geq 2\iota(x^{\prime}-x)^{2}.$ (4.31)
-Moreover, $\partial_{x}f$ is $l$-Lipschitz in $a$ and $\hat{\alpha}$ satisfies (4.22), we have that
+Moreover, $\partial_{x}f_{1}$ is $l$-Lipschitz in $a$ and $\hat{\alpha}$ satisfies (4.22), we have that
 $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))\right)$ (4.32)
@@
 $-\hat{X}\left(\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
-$\leq$ $-2\iota\hat{X}^{2}+l|\hat{X}|(\frac{l}{2\eta}|\hat{X}|+\frac{|b_{3}|}{2\eta}|\hat{Y}|).$
+[…]
 From Lemma 4.3, $\hat{\alpha}$ satisfies (4.24), it follows that
@@
 $=$ $b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X,Y^{\prime})\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y^{\prime})-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
-$\leq-\frac{2b_{3}^{2}\eta}{\zeta^{2}}\hat{Y}^{2}+|b_{3}\hat{Y}|\left(\frac{l}{2\eta}|\hat{X}|\right).$
+[…]
 By applying elementary estimates, we derive
 $\mathbb{E}\left[-r\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (4.34)
-$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2}+|b_{4}|)\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2})\mathbb{E}[\hat{Y}^{2}]$
+$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2}+2b_{4})\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2})\mathbb{E}[\hat{Y}^{2}]$
 $<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
@@
 $B^{\prime}(t,x,y)=b_{1}x+b_{2}\overline{\mu}_{t}+b_{3}\hat{\alpha}(x,y),$ (4.35)
-$F^{\prime}(t,x,y)=b_{1}y+b_{4}\overline{\mu}_{t}+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
+$F^{\prime}(t,x,y)=b_{1}y+\partial_{x}f_{0}(x,\mu_{t})+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
 Following the identical analytical procedure, we have
 $\mathbb{E}\left[-r\hat{X}\hat{Y}-\hat{X}(F^{\prime}(t,X,Y)-F^{\prime}(t,X^{\prime},Y^{\prime}))+\hat{Y}(B^{\prime}(t,X,Y)-B^{\prime}(t,X^{\prime},Y^{\prime}))\right]$ (4.36)
-$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta})\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta})\mathbb{E}[\hat{Y}^{2}]$
+$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta}+b_{4})\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta})\mathbb{E}[\hat{Y}^{2}]$
 $<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
@@
 6 Viscosity solution to distribution-dependent elliptic PDE
-6.1 Regularity of the FBSDE solutions
 We have proven that the classical solutions of the master equations can be employed to resolve the infinite-time FBSDEs.
@@
   • The value of $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$, but not on the specific realization of $\xi$.
-  • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on ×$\mathbb{R}\times\mathcal{P}_{2}$.
-While the validity of these two conditions has not been established for general cases, we can prove them for linear cases.
-If we set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, and we further assume that all assumptions in Theorem 4.4 and (4.27) are satisfied, the FBSDEs (6.1) and (6.2) become the following linear equations:
-$\begin{cases}\mathrm{d}X_{t}^{\xi}=\left(b_{1}X_{t}^{\xi}-\frac{b^{2}_{3}}{2C}Y_{t}^{\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[(b_{1}-r)Y_{t}^{\xi}+2AX_{t}^{\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.4)
-$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\left(b_{1}X_{t}^{x,\xi}-\frac{b^{2}_{3}}{2C}Y_{t}^{x,\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[(b_{1}-r)Y_{t}^{x,\xi}+2AX_{t}^{x,\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right]\mathrm{d}t+Z_{t}^{x,\xi}\mathrm{d}B_{t},\\ X_{0}^{x,\xi}=x.\end{cases}$ (6.5)
-We have already proved that for any initial state $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $x\in\mathbb{R}$, equations (6.4) and (6.5) have unique solutions in $L^{2}_{r}$.
-Lemma 6.1
-For two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution, their corresponding solutions to (6.4) have the same expectations. That is
-$\mathbb{E}[X_{t}^{\xi_{1}}]=\mathbb{E}[X_{t}^{\xi_{2}}],\quad~{}\mbox{for a.e.$t\geq 0.$}$ (6.6)
-Proof.
-Fix $\xi_{1}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and the corresponding solution $X^{\xi_{1}}_{t}$ to equation (6.4).
-Taking $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]$, we consider the following two infinite-time FBSDEs:
-$\begin{cases}\mathrm{d}X_{t}^{\xi}=\left(pX_{t}^{\xi}-qY_{t}^{\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left(uY_{t}^{\xi}+vX_{t}^{\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.7)
+  • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$.
+6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
+We shall prove the weak uniqueness for the infinite-time FBSDE:
+$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
+We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}$. We then give the definitions of strong and weak uniqueness.
+Definition 6.1 (Strong uniqueness)
+We say that the strong uniqueness holds for FBSDE (6.4) if on any admissible set-up $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
+$(X^{1}_{t},Y^{1}_{t},Z^{1}_{t})_{t\geq 0},\quad(X^{2}_{t},Y^{2}_{t},Z^{2}_{t})_{t\geq 0}$ (6.5)
+satisfying the FBSDE (6.4) with the same initial condition $X_{0}$ (up to an exceptional event), it holds that
+$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|X^{1}_{t}-X^{2}_{t}|^{2}+|Y^{1}_{t}-Y^{2}_{t}|^{2}+|Z^{1}_{t}-Z^{2}_{t}|^{2}\right)\mathrm{d}t\right]=0$ (6.6)
+Definition 6.2 (Weak uniqueness)
+For any two set-ups $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1},\mathbb{F}^{1})$ and $(\Omega^{2},\mathcal{F}^{2},\mathbb{P}^{2},\mathbb{F}^{2})$ with inputs $(X_{0}^{1},B^{1})$ and $(X_{0}^{2},B^{2})$, $X_{0}^{1},X_{0}^{2}$ have the same law on $\mathbb{R}$.
+We say the weak uniqueness holds for FBSDE (6.4) if the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on corresponding set-ups have the same distribution.
+We use the same scheme as the one developed by Yamada and Watanabe to prove that path-wise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
+Theorem 6.3
+Assume that on $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B_{t})_{t\geq 0}$, the FBSDE has a unique strong solution $(X_{t},Y_{t},Z_{t})$.
+Then the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
+Proof. At first we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
+$d(\omega^{1},\omega^{2})=\sum_{n\geq 0}2^{-n}\sup_{t\in[0,n]}\max(|\omega_{t}^{1}-\omega_{t}^{2}|,1).$ (6.7)
+And define
+$\begin{split}&\Omega_{\text{input}}\triangleq\mathbb{R}\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{output}}\triangleq C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{canon}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}.\end{split}$ (6.8)
+Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i}),i=1,2$.
+Denote by $Q^{1}$ and $Q^{2}$ the distribution of $(X_{0}^{1},B_{t}^{1},X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{0}^{2},B_{t}^{2},X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on $\Omega_{\text{canon}}=\Omega_{\text{input}}\times\Omega_{\text{output}}$, by $Q_{\text{input}}$ the common distribution of the processes $(X_{0}^{1},B_{t}^{1}),(X_{0}^{2},B_{t}^{2})$ on $\Omega_{\text{input}}$.
+Let us now define for $i\in\{1,2\}$,
+$q^{i}(\omega_{\text{input}};F):\Omega_{\text{input}}\times\mathcal{B}(\Omega_{\text{output}})\rightarrow[0,1]$ (6.9)
+as the regular conditional probability for $\mathcal{B}(\Omega_{\text{output}})$ given $\omega_{\text{input}}\in\Omega_{\text{input}}$ (under $Q^{i}$). It satisfies:
+  • $\forall\omega_{\text{input}}\in\Omega_{\text{input}}$, $q^{i}(\omega_{\text{input}};\cdot)$ is a probability measure on $(\Omega_{\text{output}},\mathcal{B}(\Omega_{\text{output}}))$.
+  • $\forall F\in\mathcal{B}(\Omega_{\text{output}})$, the mapping $(\xi,w)\rightarrow q^{i}(\xi,w;F)$ is ${\cal B}(\mathbb{R})\otimes{\cal B}(C([0,\infty),\mathbb{R}))$-measurable.
+  • $\forall F\in{\cal B}(\Omega_{\text{output}}),\forall G\in{\cal B}(\Omega_{\text{input}})$:
+    $Q^{i}(G\times F)=\int_{G}q^{i}(\omega_{\text{input}};F)Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}).$ (6.10)
+Next, we need a enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
+$\Omega_{\text{total}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}\times\Omega_{\text{output}},$ (6.11)
+and $\mathcal{G}$ is the completion of the $\sigma$-field ${\cal B}(\Omega_{\text{canon}})\otimes{\cal B}(\Omega_{\text{output}})$ by the collection $\cal N$} of all null sets under the probability measure
+$Q(G\times F_{1}\times F_{2})=\int_{G}q^{1}(\omega_{\text{input}};F_{1})q^{2}(\omega_{\text{input}};F_{2})Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}),$ (6.12)
+where $F_{1},F_{2}\in{\cal B}(\Omega_{\text{output}}),G\in{\cal B}(\Omega_{\text{input}})$.
+We observe that $Q(G\times F_{1}\times\Omega_{\text{output}})=Q^{1}(G\times F_{1}),Q(G\times\Omega_{\text{output}}\times F_{2})=Q^{2}(G\times F_{2}).$ Inparticular, we denote by $(\xi,w,x^{1},y^{1},\zeta^{1},x^{2},y^{2},\zeta^{2})$ the canonical process on $\Omega_{\text{total}}$, then $(\xi,w,x^{1},y^{1},\zeta^{1})$ has distribution $Q^{1}$ and $(\xi,w,x^{2},y^{2},\zeta^{2})$ has distribution $Q^{2}$.
+We define $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ by
+$z^{i}_{t}=\begin{cases}\lim_{n\rightarrow\infty}n\left(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+}\right)\quad~\mbox{if the limit exists,}\\ 0\quad~\mbox{otherwise.}\end{cases}$ (6.13)
+[…]
+$\zeta^{i}_{t}=\int_{0}^{t}z^{i}_{s}\mathrm{d}s,\quad t\geq 0.$ (6.14)
+Moreover, we have
+$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
+the last inequality following from Fubini’s theorem.
+Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augumentation under $Q$ of the canonical filtration
+${\cal H}_{t}=\{(\xi,w_{s},x^{1}_{s},y^{1}_{s},\zeta^{1}_{s},x^{2}_{s},y^{2}_{s},\zeta^{2}_{s});0\leq s\leq t\}$ (6.16)
+on $\Omega_{\text{total}}$.
+It’s easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable.
+Moreover, for $i\in\{1,2\}$,
+$Q\left\{\omega\in\Omega_{\text{total}};(x,w,x^{i},y^{i},\zeta^{i})\in A\right\}=Q^{i}\left\{(X_{0}^{i},B^{i},X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\right\};\quad A\in{\cal B}(\Omega_{\text{canon}}).$ (6.17)
+Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in ([6], Remark 1.6). Let us firstly define
+$\pi_{t}:C([0,\infty);\mathbb{R})\rightarrow C([0,t];\mathbb{R}),\quad h\rightarrow h|_{[0,t]},$ (6.18)
 and
-$\begin{cases}\mathrm{d}\tilde{X}_{t}^{\xi}=\left(p\tilde{X}_{t}^{\xi}-q\tilde{Y}_{t}^{\xi}+b_{2}\phi_{t}\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}\tilde{Y}_{t}^{\xi}=-\left(u\tilde{Y}_{t}^{\xi}+v\tilde{X}_{t}^{\xi}+b_{4}\phi_{t}\right)\mathrm{d}t+\tilde{Z}_{t}^{\xi}\mathrm{d}B_{t},\\ \tilde{X}_{0}^{\xi}=\xi.\end{cases}$ (6.8)
-Where $p=b_{1},q=\frac{b^{2}_{3}}{2C},u=b_{1}-r,v=2A$, and according to the argument in Theorem 4.4, each of the aforementioned equations admits a unique solution in $L^{2}_{r}$.
-We denote the solutions for Equation (6.7) with initial states ($\xi_{1},\xi_{2}$) by ($X_{t}^{\xi_{1}},X_{t}^{\xi_{2}}$), and that for Equation (6.8) under identical initial states by ($\tilde{X}_{t}^{\xi_{1}},\tilde{X}_{t}^{\xi_{2}}$).
-For Equation (6.8), let $P_{t}^{\xi}=m\tilde{X}_{t}^{\xi}-\tilde{Y}_{t}^{\xi}$ where $m$ is the positive solution of $qm^{2}-(p+u)m-v=0$. Then
-$\begin{split}\mathrm{d}P_{t}^{\xi}&=\left[(mp+v)\tilde{X}_{t}^{\xi}-(mq-u)\tilde{Y}_{t}^{\xi}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+\mathrm{d}B_{t}\\ &=\left[(mq-u)P_{t}^{\xi}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+(m-\tilde{Z}_{t}^{\xi})\mathrm{d}B_{t}.\end{split}$ (6.9)
-This demonstrates that $P^{\xi}$ satisfies an infinite-time backward stochastic differential equation
-$\mathrm{d}P_{t}=\left[(mq-u)P_{t}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+Z_{t}\mathrm{d}B_{t}$ (6.10)
-and exhibits no dependence on $\xi$.
-Applying ([14], Theorem 4) with the facts that $mq-u\geq-u=r-b_{1}\geq\frac{r}{2}$ and $\phi_{t}\in L^{2}_{r}$, the above equation admits a unique solution $(P,Z)\in L^{2}_{r}$.
-Then we know $\tilde{X}_{t}^{\xi}$ satisfy the following SDE,
-$\mathrm{d}\tilde{X}_{t}^{\xi}=\left[(p-mq)\tilde{X}_{t}^{\xi}+qP_{t}+b_{2}\phi_{t}\right]\mathrm{d}t+\mathrm{d}B_{t},\quad\quad\tilde{X}_{0}^{\xi}=\xi.$ (6.11)
-It is easy to see that $\tilde{X}_{t}^{\xi_{1}}$ and $\tilde{X}_{t}^{\xi_{2}}$ have the same distribution, so we have $\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{1}}].$
-Since $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]$, $X_{t}^{\xi_{1}}$ and $\tilde{X}_{t}^{\xi_{1}}$ are the same.
-Then we have $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{1}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]$, which means $\tilde{X}_{t}^{\xi_{2}}$ is also a solution to FBSDE (6.7) with initial $\xi_{2}$.
-By the uniqueness of the solution, we get that $\mathbb{E}[{X}_{t}^{\xi_{2}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]=\mathbb{E}[X_{t}^{\xi_{1}}]$.
-Now we finish the proof.
+$\pi^{\prime}_{t}:\gamma\rightarrow\gamma_{t},\quad h\rightarrow h|_{[0,t]},$ (6.19)
+where $\gamma=\Omega_{\text{output}}$ and
+$\gamma_{t}=C([0,t];\mathbb{R})\times C([0,t];\mathbb{R})\times C([0,t];\mathbb{R}).$ (6.20)
+Endowing $C([0,t];\mathbb{R})$ and $\gamma_{t}$ with their borelian $\sigma$-fields, we define
+${\cal K}_{t}\triangleq\sigma\{\pi_{t}\},\quad{\cal K}^{\prime}_{t}\triangleq\sigma\{\pi^{\prime}_{t}\}.$ (6.21)
+Using the separability of the spaces $C([0,t];\mathbb{R})$ and $\gamma_{t}$, we see that
+${\cal K}_{t}=\sigma\{w_{s};0\leq s\leq t\},$ (6.22)
+and that $\forall i\in\{1,2\},\forall A\in{\cal K}^{\prime}_{t}$, the set $\{(X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\}$ belongs to ${\cal F}^{i}_{t}$.
+Now, considering $A\in{\cal K}^{\prime}_{t}$ , we want to show that, for $i\in\{1,2\}$, the map
+$\Omega_{\text{input}}\rightarrow[0,1],\quad(\xi,w)\rightarrow q^{i}(\xi,w;A)$ (6.23)
+is measurable with respect to the completion of the $\sigma$-field ${\cal B}(\mathbb{R})\otimes{\cal K}_{t}$ under the probability measure $Q_{\text{input}}$, denoted $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$.
+Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$.
+Then, $\forall i\in\{1,2\}$
+$\begin{split}&\int I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)q^{i}(\xi,w;A)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\mathbb{E}^{\mathbb{P}^{i}}\left[I_{F}(X_{0})I_{G_{1}}(B^{i})I_{G_{2}}(B^{i})I_{A}(X^{i},Y^{i},\int_{0}^{\cdot}Z_{s}^{i}\mathrm{d}s)\right]\\ =&\mathbb{E}^{\mathbb{P}^{i}}\left[I_{F}(X_{0})I_{G_{1}}(B^{i})I_{A}(X^{i},Y^{i},\int_{0}^{\cdot}Z_{s}^{i}\mathrm{d}s)\right]\mathbb{E}^{\mathbb{P}^{i}}\left[I_{G_{2}}(B^{i})\right]\\ =&\int I_{F}(\xi)I_{G_{1}}(w)q^{i}(\xi,w;A)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\int I_{G_{2}}(w)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w).\end{split}$ (6.24)
+Hence, the map $(\xi,w)\rightarrow q^{i}(\xi,w;A)$ is measurable with respect to $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$.
+Now we prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion.
+Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$.
+Then,
+$\begin{split}&\mathbb{E}^{Q}\left[I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)I_{A}(x^{1},y^{1},\zeta^{1})I_{A^{\prime}}(x^{2},y^{2},\zeta^{2})\right]\\ =&\int I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)q^{1}(\xi,w;A)q^{2}(\xi,w;A^{\prime})Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\int I_{F}(\xi)I_{G_{1}}(w)q^{1}(\xi,w;A)q^{2}(\xi,w;A^{\prime})Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\int I_{G_{2}}(w)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\mathbb{E}^{Q}\left[I_{F}(\xi)I_{G_{1}}(w)I_{A}(x^{1},y^{1},\zeta^{1})I_{A^{\prime}}(x^{2},y^{2},\zeta^{2})\right]\mathbb{E}^{Q}\left[I_{G_{2}}(w)\right].\end{split}$ (6.25)
+Noting that ${\cal H}^{t}={\cal B}(\mathbb{R})\otimes{\cal K}_{t}\otimes{\cal K}^{\prime}_{t}\otimes{\cal K}^{\prime}_{t}$, we conclude that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion.
+At last, applying the same procedure on $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in ([5], Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
+$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}B(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
+Through the strong uniqueness, we know that under $Q$, the processes $(x^{1}_{t},y^{1}_{t},\zeta^{1}_{t})_{t\geq 0}$ and $(x^{2}_{t},y^{2}_{t},\zeta^{2}_{t})_{t\geq 0}$ have the same law, then we get the desired result.
 Remark 6.4
 For any two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution and $x\in\mathbb{R}$, the solutions $Y^{x,\xi_{1}},Y^{x,\xi_{2}}$ to (6.2) are the same.
-We can say that $Y^{x,\xi}_{0}$ depends depends solely on the distribution of $\xi$.
+We can say that $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$.
+6.2 Regularity of the FBSDE solutions
 Lemma 6.5
 For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$.
-Then we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
+Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
 $|\mathcal{V}(x,\mu)-\mathcal{V}(x^{\prime},\mu^{\prime})\leq C(|x-x^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime}))$ (6.27)
@@
 Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{x_{1},\xi_{1}}_{t}-Y^{x_{2},\xi_{2}}_{t}|^{2}$, we get
-$\begin{split}|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}&\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\mathbb{E}\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t\\ &\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t.\end{split}$ (6.14)
+$\begin{split}|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}&\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+W_{2}^{2}(X^{\xi_{1}}_{t},X^{\xi_{2}}_{t})\right]\mathrm{d}t\\ &\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t.\end{split}$ (6.29)
 Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{x,\xi}\hat{Y}^{x,\xi}$, we get
@@
 Assume that both FBSDEs (6.1) and (6.2) admit unique solutions in $L^{2}_{r}$, $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$.
-Then we define the function $\mathcal{V}(x,\mu)\triangleq Y_{0}^{x,\xi}$ with $\mathcal{L}_{\xi}=\mu$, and assume $\mathcal{V}$ is jointly continuous on ×$\mathbb{R}\times\mathcal{P}_{2}$.
+Then we define the function $\mathcal{V}(x,\mu)\triangleq Y_{0}^{x,\xi}$ with $\mathcal{L}_{\xi}=\mu$, and assume $\mathcal{V}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$.
 The function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.36).

### v4

@@
 This is precisely the relationship expressed in Equation (1.4).
+Alternatively, if we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$, we demonstrate that $\mathcal{V}$ serves as a viscosity solution to the distribution-dependent PDE below:
+$\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (1.11)
 This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we prove the Yamada-Watanabe theorem for infinite-time FBSDEs and provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
@@
 To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$. For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
-$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
+$\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
 It has a unique solution $(X_{t},Y_{t},Z_{t})\in L^{2}_{K}(0,+\infty,\mathbb{R}^{3})$ under the following assumptions:
@@
 (i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
-$|B(t,x,y,\mu)$ $-B(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
+$|G(t,x,y,\mu)$ $-G(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
 $\leq\ell(|x-x^{\prime}|+|y-y^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime})).\quad\mathrm{a.s.}$
@@
 Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5].
-For $i=1,2$, let $dX^{i}_{t}\triangleq b^{i}_{t}dt+\sigma^{i}_{t}dB_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}:={\cal L}_{X^{2}_{t}}$.
+For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$.
 Fix $T>0$ and let all conditions be restricted to the interval $[0,T]$.
 Then we have
-$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big{[}\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\hbox{\rm tr$\,$}\big{(}\partial_{xx}U[\sigma_{t}^{1}]^{2}\big{)}\Big{]}(t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big{(}\tilde{\mathbb{E}}_{{\cal F}_{t}}\big{[}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big{]}\Big{)}\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
+$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U(\sigma_{t}^{1})^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{\sigma}_{t}^{2})^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
 Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ .
@@
 $\hat{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))=(b_{1}-\frac{a_{1}b_{3}^{2}}{C})x+(b_{2}-\frac{a_{2}b_{3}^{2}}{2C})\bar{\mu}.$ (5.20)
-Since $b_{1}+|b_{2}|<\frac{r}{2}$, if we further require $a_{1}>0,a_{2}>0$, the SDE (5.3) admits a unique solution in $L_{r}^{2}$.
+Since the SDE (5.3) now becomes linear, if we further require $a_{1}>0,a_{2}>0$, it’s not hard to see that the SDE admits a unique solution in $L_{r}^{2}$.
 6 Viscosity solution to distribution-dependent elliptic PDE
@@
 We shall prove the weak uniqueness for the infinite-time FBSDE:
-$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
-We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}$. We then give the definitions of strong and weak uniqueness.
+$\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
+We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}\in\mathbb{L}^{2}({\cal F}_{0})$.
+We then give the definitions of strong and weak uniqueness.
 Definition 6.1 (Strong uniqueness)
@@
 $\begin{split}&\Omega_{\text{input}}\triangleq\mathbb{R}\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{output}}\triangleq C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{canon}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}.\end{split}$ (6.8)
-Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i}),i=1,2$.
+Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i})$ with identically distributed inputs $(X_{0}^{i},B^{i})$, $i=1,2$, on each of which a solution $(X_{t}^{i},Y_{t}^{i},\int_{0}^{t}Z^{i}_{s}\mathrm{d}s)_{t\geq 0}$ to the FBSDE (6.4) is defined.
 Denote by $Q^{1}$ and $Q^{2}$ the distribution of $(X_{0}^{1},B_{t}^{1},X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{0}^{2},B_{t}^{2},X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on $\Omega_{\text{canon}}=\Omega_{\text{input}}\times\Omega_{\text{output}}$, by $Q_{\text{input}}$ the common distribution of the processes $(X_{0}^{1},B_{t}^{1}),(X_{0}^{2},B_{t}^{2})$ on $\Omega_{\text{input}}$.
@@
 Moreover, we have
-$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
+$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{-rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{-rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{-rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{-rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{-rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
 the last inequality following from Fubini’s theorem.
@@
 is measurable with respect to the completion of the $\sigma$-field ${\cal B}(\mathbb{R})\otimes{\cal K}_{t}$ under the probability measure $Q_{\text{input}}$, denoted $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$.
-Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$.
+Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\sigma\{w_{s}-w_{t};s\geq t\}$.
 Then, $\forall i\in\{1,2\}$
@@
 Now we prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion.
-Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$.
+Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\sigma\{w_{s}-w_{t};s\geq t\}$.
 Then,
@@
 At last, applying the same procedure on $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in ([5], Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
-$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}B(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
+$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}G(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
 Through the strong uniqueness, we know that under $Q$, the processes $(x^{1}_{t},y^{1}_{t},\zeta^{1}_{t})_{t\geq 0}$ and $(x^{2}_{t},y^{2}_{t},\zeta^{2}_{t})_{t\geq 0}$ have the same law, then we get the desired result.
@@
 Now let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
-$(\mathcal{L}\mathcal{U})(x,\mu,\mathcal{U}(x,\mu))+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.36)
+$(\mathcal{L}\mathcal{U})[\mathcal{U}](x,\mu)+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.36)
 where
-$\begin{split}(\mathcal{L}\Phi)(x,\mu,\Psi)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
+$\begin{split}(\mathcal{L}\Phi)[\Psi](x,\mu)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
 and
-$F(x,\mu,\Psi)\triangleq\partial_{x}H(x,\mu,\Psi(x,\mu))-r\Psi(x,\mu).$ (6.38)
+$F(x,\mu,\Psi(x,\mu))\triangleq\partial_{x}H(x,\mu,\Psi(x,\mu))-r\Psi(x,\mu).$ (6.38)
+Definition 6.6
+A function $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ is said to be a test function if the quantities:
+[…]
+and
+[…]
+are finite, uniformly in $(x,\mu)$ in any compact subset of $\mathbb{R}\times\mathcal{P}_{2}$.
 Definition 6.7
@@
 Then $\mathcal{U}$ is called a viscosity subsolution (resp.
-supersolution) of PDE (6.36) if, whenever $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp.
+supersolution) of PDE (6.36) if, whenever $\Psi$ is a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp.
 minimum) of $\mathcal{U}-\Psi$, we have
-$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.39)
+$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.41)
 (respectively
-$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.40)
-). The function $\mathcal{U}$ is called a viscosity solution of PDE (6.36) if it is both a viscosity subsolution and a viscosity supersolution.
+$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.42)
+The function $\mathcal{U}$ is called a viscosity solution of PDE (6.36) if it is both a viscosity subsolution and a viscosity supersolution.
 We now assert that
@@
 Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$.
-Let $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum of $\mathcal{U}-\Psi$.
+Let $\Psi$ be a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $\mathcal{U}-\Psi$.
 We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$.
 We suppose that
-$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.41)
+$(\mathcal{L}\Psi)[\mathcal{V}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.43)
 It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
-$\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)(x,\mu,\mathcal{V}(x,\mu))+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.42)
-Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2).
+$\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)[\mathcal{V}](x,\mu)+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.44)
+Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu^{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2).
 For some $T>0$, let $\tau$ denote the stopping time
-$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\mathcal{L}_{X_{t}^{\xi^{0}}})\in O\}\land T.$ (6.43)
-We first note that the pair of processes
+$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\mathcal{L}_{X_{t}^{\xi^{0}}})\notin O\}\land T.$ (6.45)
+Let $\rho_{t}\triangleq{\cal L}_{X_{t}^{\xi^{0}}}$, we first note that the pair of processes
 $(\bar{Y}\left(t\right),\bar{Z}\left(t\right))\triangleq(Y^{x^{0},\xi^{0}}_{t\wedge\tau},I_{\left[0,\tau\right]}(t)Z^{x^{0},\xi^{0}}_{t}),\quad 0\leq t\leq T,$ (6.46)
 is the solution of BSDE
-$\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.45)
+$\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\rho_{s},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\rho_{s}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.47)
 Next, it follows from Itô’s formula that the pair of processes
-$(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}))$ (6.46)
+$(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\rho_{t\wedge\tau}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\rho_{t\wedge\tau}))$ (6.48)
 is the solution of BSDE
-$\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.47)
+$\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi[\mathcal{V}](X^{x^{0},\xi^{0}}_{s},\rho_{s})\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.49)
 Define
-$\beta_{s}=-\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))-F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))$ (6.48)
+$\beta_{s}=-\mathcal{L}\Psi[\mathcal{V}](X^{x^{0},\xi^{0}}_{s},\rho_{s})-F(X^{x^{0},\xi^{0}}_{s},\rho_{s},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\rho_{s}))$ (6.50)
 and
@@
 We have
-$\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.50)
+$\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.52)
 Therefore
@@
 Now from the choice of $O$ and $\tau$, a.s.
-$\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.52)
+$\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.54)
 Consequently, $\tilde{Y}_{0}=\Psi(x^{0},\mu^{0})-\mathcal{V}(x^{0},\mu^{0})>0$, which contradicts the earlier assumption.

### v5

@@
 To solve this problem, we partition all agents into a representative player and the social equilibrium.
-When the optimal strategy of the representative player shares the same feedback form with the strategy of the social equilibrium, we say the system achieves a Nash equilibrium.
+When the optimal strategy of the representative player has the same feedback form as the strategy in the social equilibrium, we say that the system achieves a Nash equilibrium.
 We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations (FBSDEs).
-By employing the elliptic master equations, a class of distribution-dependent elliptic PDEs , we provide a representation for the Nash equilibrium.
-We prove the Yamada-Watanabe theorem and show the weak uniqueness for infinite-time FBSDEs.
-And we prove that the solutions to a system of infinite-time FBSDEs can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
+By employing elliptic master equations, a class of distribution-dependent elliptic partial differential equations (PDEs), we provide a representation for the Nash equilibrium strategies.
+We prove the Yamada-Watanabe type theorem and show weak uniqueness for infinite-time FBSDEs.
+Furthermore, we prove that the solutions to a system of infinite-time FBSDEs can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
 Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, weak uniqueness, elliptic master equations.
 1 Introduction
-The study of mean field games was initiated independently by Lasry-Lions (see [9, 10, 11]) and Huang-Malhamé-Caines [8], which is an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games.
-It is noteworthy that current theoretical frameworks are primarily developed for finite-time problems, while infinite-time scenarios remain significantly underdeveloped.
+The study of mean field games was initiated independently by Lasry-Lions [11, 12, 13] and Huang-Malhamé-Caines [10] as an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games.
+It is noteworthy that current theoretical frameworks are primarily developed for finite-time problems, while infinite-time scenarios remain considerably underdeveloped.
 We refer the reader to [5] for a comprehensive exposition on this subject.
-In this paper, we consider a generalized framework for mean field games, which extends the classical finite-time settings to discounted infinite-time mean field games.
-Our model differs from the fixed-point problems studied in prior works.
+In this paper, we consider a generalized framework for mean field games, which extends classical finite-time settings to discounted infinite-time mean field games.
 In our framework, a representative player interacts with a continuum of other players (also referred to as the population or social equilibrium).
-The dynamic of the private state of the representative player is given by
-$X^{x}_{t}=x+\int_{0}^{t}b(X^{x}_{t},\mu_{s},\beta_{s}^{x})\mathrm{d}s+B_{t}.$ (1.1)
-Here $\mu_{t}$ is the population distribution , while $\beta^{x}$ is the optimal strategy of the representative player, derived as the solution to the stochastic optimization problem:
-$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right],$ (1.2)
+Let $\mu_{0}\in{\cal P}(\mathbb{R})$ be the initial state distribution in the society, and let $(\mu_{t})_{t\geq 0}$ denote the population distribution flow starting at $\mu_{0}$.
+For the player($x$), a representative player with initial state $x$, the dynamic of its state is given by
+$X^{x,\beta}_{t}=x+\int_{0}^{t}b(X^{x,\beta}_{t},\mu_{s},\beta_{s})\mathrm{d}s+B_{t},$ (1.1)
+where $\beta$, an $\mathbb{R}$-valued progressively measurable stochastic process, is the strategy of the player($x$).
+Denote by $\beta^{x}$ the optimal strategy of the representative player($x$) derived from the stochastic optimization problem:
+$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (1.2)
+Since the population consists of a multitude of individuals, the macroscopic distribution should satisfy
+$\mu_{t}(\cdot)=\int\mathbb{P}(X_{t}^{x,\beta^{x}}\in\cdot)\mu_{0}(\mathrm{d}x).$ (1.3)
+This is exactly a fixed-point problem.
+Note that our definition of mean field games differs from that in [1] and from those of earlier finite-time mean field games.
+In [1], all agents are identical and thus represented by a single representative player.
+For a given measure flow $(\mu_{t})_{t\geq 0}$, the representative player seeks to minimize
+$J^{\mu}(\alpha)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(t,X_{t},\mu_{t},\alpha_{t})\mathrm{d}t\right]$ (1.4)
 under the constraint
-$\begin{cases}\mathrm{d}X_{t}^{x,\beta}=b(X_{t}^{x,\beta},\mu_{t},{\beta}_{t})\mathrm{d}t+\mathrm{d}B_{t},\\ X_{0}^{x,\beta}=x.\end{cases}$ (1.3)
-Since the population consists of a multitude of homogeneous individuals, the macroscopic state should satisfy
-$X^{\xi}_{t}=X^{x,\beta^{x}}|_{x=\xi},$ (1.4)
-where $\mathcal{L}_{\xi}=\mu_{0}$ denotes the initial population distribution. We say the population reaches equilibrium if $\mu_{t}=\mathcal{L}_{X^{\xi}_{t}}.$
-This model decouples the micro-level agent from the macro-level societal distribution, enabling interconnected analysis of their evolution.
-The equilibrium is characterized by two consistency conditions:
-  • Individual Rationality: The representative player’s optimal strategy is consistent with the perceived social equilibrium.
-  • Macro Consistency: The aggregate distribution generated by all players adopting this strategy must equal the posited social equilibrium.
-We will employ the stochastic maximum principle and infinite-time forward-backward stochastic differential equations (FBSDEs) to solve this game-theoretic problem, and derive a representation of the equilibrium strategy via an elliptic-type master equation.
-The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [13, 14] in the early 1990s, which is now a typical tool in stochastic optimization problems.
-We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time forward-backward stochastic differential equations.
-The foundational work in [15] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, and subsequent work in [16] investigated a more general setting and established connections with quasilinear elliptic PDEs.
-Recently, [1] extended this framework to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games.
-In this paper, we partition all agents into a representative player and the social equilibrium, and characterize the equilibrium state through the following system of infinite-time FBSDE:
+$\begin{cases}\mathrm{d}X_{t}=b\left(t,X_{t},\mu_{t},{\alpha}_{t}\right)\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (1.5)
+Then, we require that the law of $X_{t}$ coincides with $\mu_{t}$, which means that we need to find a fixed point.
+In the conventional definition of mean field games, the primary focus on the distribution flow $(\mu_{t})_{t\geq 0}$ stems from its role as the limit of the Nash equilibrium of an $N$-player game.
+Within this framework, the master equation also serves as a crucial tool: see [7] and [8].
+In the study by Buckdahn et al.
+[2], the individual state process and the population state process are treated separately, allowing for an investigation into various properties of the parabolic master equation.
+Inspired by their approach, our definition of mean field games explicitly decouples the representative player from the population.
+This separation offers two key advantages: first, it provides a mathematically interpretation of the game problem even in the absence of a realistic $N$-player game context; second, the value function $V(x,\mu_{0})$ of the representative player($x$) directly embodies the relationship between the individual state $x$ and the initial population distribution $\mu_{0}$, thereby laying a solid foundation for further analysis of the value function’s properties.
+We employ the stochastic maximum principle and infinite-time forward-backward stochastic differential equations (FBSDEs) to solve this game-theoretic problem and derive a representation of the equilibrium strategy via an elliptic master equation.
+The theory of general nonlinear BSDEs was pioneered by Pardoux and Peng [15, 16] in the early 1990s, which is now a typical tool in stochastic optimization problems.
+We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time FBSDEs.
+The foundational work in [17] proved the existence and uniqueness of solutions to infinite-time FBSDEs, and subsequent work in [18] investigated a more general setting and established connections with quasilinear elliptic partial differential equations (PDEs).
+Recently, [1] extended this framework to McKean-Vlasov FBSDEs, which play crucial roles in obtaining equilibrium solutions for mean field games.
+In this paper, we partition all agents into a representative player and the social equilibrium and characterize the equilibrium state through the following system of infinite-time FBSDE:
 $\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.6)
 $\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.7)
-Here $\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy$ is the generalized Hamiltonian, and $\hat{\alpha}(x,y)$ is the minimizer of $\mathcal{H}(x,\mu,a,y)$ with respect to $a$ when we assume $\mathcal{H}$ is separable in variables $\mu$ and $a$.
+Here, $\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy$ is the generalized Hamiltonian, and $\hat{\alpha}(x,y)$ is the minimizer of $\mathcal{H}(x,\mu,a,y)$ with respect to $a$ when we assume that $\mathcal{H}$ is separable in variables $\mu$ and $a$.
+It is worth emphasizing that we set the diffusion coefficient to 1 here, which makes both equations (1.6) and (1.7) solvable.
+If the diffusion term in problem setup (1.1) is dependent on the state, distribution, and control, studying these two equations would be considerably more challenging.
+Now, we consider the solutions to these two FBSDEs.
 The solution $X_{t}^{\xi}$ to Equation (1.6) represents the population’s state process, whose law corresponds to the population distribution $\mu_{t}$.
-And the solution $X_{t}^{x}$ to Equation (1.6) is the state process of the representative agent after solving the optimization problem.
+In addition, the solution $X_{t}^{x}$ to Equation (1.7) is the state process of the representative agent after solving the optimization problem.
 Notably, it exhibits the same feedback structure as the population’s state process.
-To further elucidate the relationship in Equation (1.4 ), we introduce the elliptic-type master equations.
-First introduced by Lions in lectures [12], the parabolic-type master equations appeared in the context of the theory of mean field games.
+To further elucidate the relationship in Equation (1.3), we introduce elliptic master equations.
+First introduced by Lions in lectures [14], the parabolic master equation appeared in the context of the theory of mean field games.
 This is a time-dependent equation that bears profound connections with finite-time mean field game theory.
@@
 We refer the reader to [5, 3, 9] for a comprehensive exposition on the subject.
-In this paper, we propose the elliptic-type master equations, which explicitly characterizes the feedback forms of both the representative player and the social equilibrium.
-This equation takes the following form:
+In this paper, we propose elliptic master equations, which explicitly characterize the feedback forms of both the representative player and the social equilibrium.
+These equations take the following form:
 $\begin{split}rU(x,\mu)=&H(x,\mu,\partial_{x}U(x,\mu))+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (1.8)
-Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are $W_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law.
-Under the assumption that the master equation (1.7) admits a solution with sufficient regularity, we derive the following representation for Equation (1.5) and (1.6):
+Here, $\partial_{x}$ and $\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu}$ and $\partial_{\tilde{x}\mu}$ are $W_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law.
+Under the assumption that the master equation (1.8) admits a solution with sufficient regularity, we derive the following representation for Equations (1.6) and (1.7):
 $Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (1.9)
 $Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (1.10)
-If we further assume that $\tilde{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\mu,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any fixed finite time $T$, we have
-$X_{t}^{x}|_{x=\xi}=X_{t}^{\xi},\quad t\in[0,T].$ (1.10)
-This is precisely the relationship expressed in Equation (1.4).
+If $\tilde{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$, we have
+$X_{t}^{x}|_{x=\xi}=X_{t}^{\xi},\quad t\in[0,T]$ (1.11)
+for any fixed finite time $T$. This is precisely the relationship expressed in Equation (1.3).
 Alternatively, if we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$, we demonstrate that $\mathcal{V}$ serves as a viscosity solution to the distribution-dependent PDE below:
 $\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (1.12)
-This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we prove the Yamada-Watanabe theorem for infinite-time FBSDEs and provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
+This paper is organized as follows: in section 2, we present preliminaries for problems in this paper; in section 3, we introduce infinite-time mean field games with discounting and the definition of a Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs; in section 5, we introduce an elliptic master equation to provide a representation for the Nash equilibrium; and in section 6, we prove the Yamada-Watanabe theorem for infinite-time FBSDEs and provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
 2 Preliminaries
-We will use the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$.
-Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any measure on $\mathbb{R}$ with finite second moment.
-Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$, define the larger filtered probability space by
+We will use a filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$.
+Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$ such that it can support any measure on $\mathbb{R}$ with finite second moment.
+Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$.
+We define the larger filtered probability space by
 $\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}$.
@@
 Let ${\cal P}\triangleq{\cal P}(\mathbb{R})$ be the set of all probability measures on $\mathbb{R}$ and let ${\cal P}_{p}(p\geq 1)$ denote the set of $\mu\in{\cal P}$ with finite $p$-th moment.
-For any sub-$\sigma$-field $\cal G$$\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, $\cal G$}-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ .
-For any $\mu,\nu\in{\cal P}_{p}$, we define the $W_{p}$–Wasserstein distance between them as follows:
-$W_{p}(\mu,\nu):=\inf\Big\{\big(\mathbb{E}[|\xi-\eta|^{q}]\big)^{1/q}:~\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big\}.$
-Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}$-valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
+For any sub-$\sigma$-field $\cal G$$\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, $\cal G$}-measurable, and $p$-integrable random variables $\xi$ and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ .
+For any $\mu,\nu\in{\cal P}_{p}$, we define the ${\cal W}_{p}$–Wasserstein distance between them as follows:
+${\cal W}_{p}(\mu,\nu)\triangleq\inf\Big\{\big(\mathbb{E}[|\xi-\eta|^{q}]\big)^{1/q}:\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big\}.$
+Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}$-valued adapted stochastic processes ($v_{t}$) starting from $t_{0}$ such that
 $\mathbb{E}\left[\int_{t_{0}}^{\infty}e^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
-To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$. For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
+To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$. For each $\mathcal{F}_{0}$-measurable square-integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
 $\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
-It has a unique solution $(X_{t},Y_{t},Z_{t})\in L^{2}_{K}(0,+\infty,\mathbb{R}^{3})$ under the following assumptions:
+Here, $G,F:\mathbb{R}^{+}\times\mathbb{R}^{2}\times{\cal P}_{2}\to\mathbb{R}$ are two measurable functions that satisfy the following assumptions.
 Assumption 2.1
-(i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
+(i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in{\cal P}_{2}$,
 $|G(t,x,y,\mu)$ $-G(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
 $\leq\ell(|x-x^{\prime}|+|y-y^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime})).\quad\mathrm{a.s.}$
-(ii) There exist constants $0<K<2\kappa$ such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$
+(ii) There exist constants $0<K<2\kappa$ such that for any $t\geq 0$ and any square-integrable random variables $X,X^{\prime},Y,Y^{\prime}$,
 $\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (2.5)
@@
 where $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
-For the detailed proof, we refer the reader to ([1], Theorem 2.1).
-We introduce the Wasserstein space and differential calculus on Wasserstein space.
-For a $W_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$, its $W_{2}$-Wasserstein derivatives[5](also called Lions-derivative), takes the form $\partial_{\mu}U:(\mu,\tilde{x})\in{\cal P}_{2}\times\mathbb{R}\to\mathbb{R}$ and satisfies:
+The following lemma establishes the existence and uniqueness of a solution to the FBSDE(2.3). For a detailed proof, we refer the reader to [1] (Theorem 2.1).
+Lemma 2.2
+Under Assumption 2.1, the FBSDE (2.3) admits a unique solution in $L_{K}^{2}$.
+We introduce the Wasserstein space and the associated differential calculus.
+For a ${\cal W}_{2}$-continuous function $U:{\cal P}_{2}\to\mathbb{R}$, its ${\cal W}_{2}$-Wasserstein derivative[5](also called the Lions-derivative) takes the form $\partial_{\mu}U:(\mu,\tilde{x})\in{\cal P}_{2}\times\mathbb{R}\to\mathbb{R}$ and satisfies
 $U({\cal L}_{\xi+\eta})-U(\mu)=\mathbb{E}\big[\langle\partial_{\mu}U(\mu,\xi),\eta\rangle\big]+o(\|\eta\|_{2}),\ \forall\ \xi\in\mathbb{L}^{2}(\mathcal{F};\mu),\eta\in\mathbb{L}^{2}(\mathcal{F}).$ (2.6)
-Let ${\cal C}^{0}({\cal P}_{2})$ denote the set of $W_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$.
-For ${\cal C}^{1}({\cal P}_{2})$, we mean the space of functions $U\in{\cal C}^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.6).
-Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
-Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}^{d}\times{\cal P}_{2}\times\mathbb{R}$.
-Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5].
-For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$.
-Fix $T>0$ and let all conditions be restricted to the interval $[0,T]$.
-Then we have
+Let $C^{0}({\cal P}_{2})$ denote the set of ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$.
+For $C^{1}({\cal P}_{2})$, we define the space of functions $U\in C^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists, is continuous on ${\cal P}_{2}\times\mathbb{R}$, and is uniquely determined by (2.6).
+Let $C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$ and $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
+Finally, we consider the space $\Theta\triangleq[0,T]\times\mathbb{R}\times{\cal P}_{2}$ for some $T>0$; let $C^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ that has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in C^{1,2,1}(\Theta)$ is that they satisfy the Itô’s formula[2, 5].
+For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,T]\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,T]\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and let $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$.
+Suppose that for every compact subset ${\cal K}\subset\mathbb{R}\times{\cal P}_{2}$, it holds that
+[…]
+We have
 $\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U(\sigma_{t}^{1})^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{\sigma}_{t}^{2})^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.8)
-Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ .
+Here, $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations under $\tilde{\mathbb{P}}$ given ${\cal F}_{t}$ and the process $(\tilde{X}^{2}_{t},\tilde{b}^{2}_{t},\tilde{\sigma}_{t}^{2})_{0\leq t\leq T}$ is a copy of the process $(X^{2}_{t},b^{2}_{t},\sigma_{t}^{2})_{0\leq t\leq T}$, defined on a copy $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ of the probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$.
 3 Infinite-time mean field games with discounting
-In this section, we introduce the infinite-time mean field games with discounting.
+In this section, we introduce infinite-time mean field games with discounting.
 Let $r>0$ represent the time discount factor and $A\subset\mathbb{R}$ be a convex control space.
-Define $\mathcal{A}\triangleq L^{2}_{r}(0,\infty,A)$ to be the space of all adimissible controls, and $b,f:\mathbb{R}\times{\cal P}_{2}\times A\rightarrow\mathbb{R}$ are two measurable functions.
+Define $\mathcal{A}\triangleq L^{2}_{r}(0,\infty,A)$ to be the space of all admissible controls, and $b,f:\mathbb{R}\times{\cal P}_{2}\times A\rightarrow\mathbb{R}$ are two measurable functions.
 We consider a population consisting of a continuum of players, where each individual player strategically interacts with the aggregate distribution formed by all other players to minimize their own cost.
@@
 $X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mu_{t},\beta_{s})\mathrm{d}s+B_{t},$ (3.1)
-where $\beta\in\mathcal{A}$ is the strategy remains to be determined. The representative player seeks to minimize the cost
+where $\beta\in\mathcal{A}$ is the strategy that remains to be determined. The representative player seeks to minimize the cost
 $J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (3.2)
-Assuming $\beta^{x}\in\mathcal{A}$ minimizes the cost function, the state process of the representative player is $X^{x,\beta^{x}}_{t}$.
-Since the representative agent can characterize the strategies of all players in the population, we assert that the following fundamental relationship must hold:
+Assuming that $\beta^{x}\in\mathcal{A}$ minimizes the cost function, the state process of the representative player is $X^{x,\beta^{x}}_{t}$.
+Since the representative player can characterize the strategies of all players with the same initial state in the population, we assert that the following fundamental relationship must hold:
 $\mu_{t}=\mathcal{L}_{X^{x,\beta^{x}}_{t}|_{x=\xi}}.$ (3.3)
-To solve this mean field game problem, we work under the following assumptions:
+To solve this mean field game problem, we work under the following assumptions.
 Assumption 3.1
@@
 $(x-x^{\prime})\left(b(x,\mu,a)-b(x^{\prime},\mu,a)\right)\leq-\lambda(x-x^{\prime})^{2}.$ (3.5)
-These assumptions jointly ensure that: (i) the state process remains confined within the $L^{2}_{r}$ space, and (ii) the cost functional maintains integrability.
-Then we partition all agents into two components: (i) the representative player, who dynamically optimizes their strategy, and (ii) the social equilibrium (or mean field), characterizing the macroscopic state shared by the population.
-For the social equilibrium, its state is governed by the following SDE:
+These assumptions jointly ensure that (i) the state process remains confined within the $L^{2}_{r}$ space and (ii) the cost functional maintains integrability.
+Then, we partition all agents into two components: (i) the representative player, who dynamically optimizes their strategy, and (ii) the social equilibrium (or mean field), characterizing the macroscopic state shared by the population.
+The state of the social equilibrium is governed by the following stochastic differential equation (SDE):
 $X_{t}^{\xi,\alpha}=\xi+\int_{0}^{t}b(X_{s}^{\xi,\alpha},\mathcal{L}_{X_{s}^{\xi,\alpha}},\alpha_{s})\mathrm{d}s+B_{t},$ (3.6)
 where $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$.
-We note that, by assumption 3.1, this SDE has a unique strong solution in $L^{2}_{r}$, see ([1], Proposition 2.2) for more details.
+We note that by assumption 3.1, this SDE has a unique strong solution in $L^{2}_{r}$; see [1] (Proposition 2.2) for more details.
 The state of the representative player is governed by
@@
 Definition 3.2
-We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
+We say that a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ constitutes a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
 $\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\xi_{0};\alpha^{\xi_{0}},\beta).$ (3.10)
-When this Nash equilibrium $\alpha^{*}$ exists, we have
+When the Nash equilibrium $\alpha^{*}$ exists, we have
 $\xi,\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad\mathcal{L}_{X_{t}^{\xi,\alpha^{\xi}}}=\mathcal{L}_{X_{t}^{\xi^{\prime},\alpha^{\xi^{\prime}}}},\quad~\mbox{for a.e.
@@
 $V(x,\mu)\triangleq J(x,\xi_{0};\alpha^{\xi_{0}},\alpha^{x}),\quad\quad\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.12)
-Note that our definition of Nash equilibrium differs from that in [1] and earlier finite-time mean field games.
-In [1], the mean field game is formulated as a fixed-point problem.
-The population is assumed to be homogeneous, meaning all agents are identical and thus represented by a single representative player.
-For a given measure flow $(\mu_{t})_{t\geq 0}$ , the representative player wants to minimize
-$J^{\mu}(\alpha)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(t,X_{t},\mu_{t},\alpha_{t})\mathrm{d}t\right],$ (3.13)
-under the constraint
-$\begin{cases}\mathrm{d}X_{t}=b\left(t,X_{t},\mu_{t},{\alpha}_{t}\right)\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (3.14)
-Then we require that the law of $X_{t}$ coincides with $\mu_{t}$, which means we need to find a fixed point.
-In this paper, we separate a representative player from the population, who only needs to consider their own optimization problem starting from state $x$.
-The representative player’s state evolution depends on both their own state $X^{x,\beta}$ and the overall population distribution $\mathcal{L}_{X^{\xi,\alpha}}$.
-Here we would like to emphasize that since there are a large number of players, any change of a representative player doesn’t impact the measure flow $\mathcal{L}_{X^{\xi,\alpha}}$.
-Under the existence assumption of the Nash equilibrium $\alpha^{*}$ specified in Definition 3.2, the stochastic dynamics of both the population and representative player are characterized by the following stochastic differential equations:
+In our framework, we separate a representative player from the population, who only needs to consider their optimization problem starting from state $x$.
+This model decouples the micro-level agent from the macro-level population distribution, enabling an interconnected analysis of their evolution.
+The equilibrium is characterized by two consistency conditions.
+  • Individual Rationality: The representative player’s optimal strategy is consistent with the perceived social equilibrium.
+  • Macro-consistency: The aggregate distribution generated by all players adopting this strategy must coincide with the posited social equilibrium.
+The representative player’s state evolution depends on both their state $X^{x,\beta}$ and the overall population distribution $\mathcal{L}_{X^{\xi,\alpha}}$.
+Here, we would like to emphasize that since the number of players is large, any change by a representative player does not impact the measure flow $\mathcal{L}_{X^{\xi,\alpha}}$.
+Under the existence assumption of the Nash equilibrium $\alpha^{*}$ specified in Definition 3.2, the stochastic dynamics of both the population and representative player are characterized by the following SDE:
 $\begin{cases}X_{t}^{\xi,\alpha^{*}}=\xi+\int_{0}^{t}b(X_{s}^{\xi,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}},{\alpha}^{*}(X_{s}^{\xi,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}}))\mathrm{d}s+B_{t},\\ X_{t}^{x,\alpha^{*}}=x+\int_{0}^{t}b(X_{s}^{x,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}},{\alpha}^{*}(X_{s}^{x,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}}))\mathrm{d}s+B_{t}.\end{cases}$ (3.13)
@@
 4 Connection with infinite-time McKean-Vlasov FBSDEs
-In this section, we employ the maximum principle to solve the optimization problem for the representative player and then use the infinite-time McKean-Vlasov FBSDE to construct the optimal strategy for the representative player such that the controls of the representative player and the social equilibrium share the same feedback form.
+In this section, we employ the maximum principle to solve the optimization problem for the representative player and then use infinite-time McKean-Vlasov FBSDEs to construct the optimal strategy for the representative player such that the controls of the representative player and the social equilibrium share the same feedback form.
 Our derivation is based on the following key assumptions on $b,f$:
 Assumption 4.1
-(i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$.
-Where $b_{0},f_{0}$ are measurable functions on $\mathbb{R}\times{\cal P}_{2}$, $b_{1},f_{1}$ are measurable functions on $\mathbb{R}\times A$.
-(ii) $b,f$ are differentiable with respect to $(x,a)$ and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
+(i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$ where $b_{0},f_{0}$ are measurable functions on $\mathbb{R}\times{\cal P}_{2}$ and $b_{1},f_{1}$ are measurable functions on $\mathbb{R}\times A$.
+(ii) $b,f$ are differentiable with respect to $(x,a)$, and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
 (iii) $H_{0}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)$ is convex with respect to $(x,a)$.
-$\min_{a\in A}H_{0}(x,\mu,a,y)$ has a unique minimizer $\hat{\alpha}(x,y)$ which is Lipschitz continuous in $(x,y)$.
+$\min_{a\in A}H_{0}(x,\mu,a,y)$ has a unique minimizer $\hat{\alpha}(x,y)$ that is Lipschitz continuous in $(x,y)$.
 4.1 Pontryagin’s maximum principle
-Assuming the state of the social equilibrium $X_{t}^{\xi}$ is given, we consider the optimization problem for the representative player, whose state is given by
+Assuming that the state of the social equilibrium $X_{t}^{\xi}$ is given, we consider the optimization problem for the representative player, whose state is given by
 $X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi}},\beta_{s})\mathrm{d}s+B_{t}.$ (4.1)
@@
 Let us formally derive the maximum principle for the infinite-time control problem.
-Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ .
+Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, and denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ .
 Let
 $R_{t}=\lim_{\epsilon\to 0}\frac{X_{t}^{x,\beta+\epsilon\gamma}-X_{t}^{x,\beta}}{\epsilon}$ (4.4)
-be the variation process. Then it can be shown that $R$ satisfies
+be the variation process. Then, it can be shown that $R$ satisfies
 $\begin{cases}\mathrm{d}R_{t}&=\left(\partial_{x}b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot R_{t}+\partial_{a}b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot\gamma_{t}\right)\mathrm{d}t,\\ R_{0}&=0.\end{cases}$ (4.5)
@@
 $\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy,$ (4.7)
-and introduce the adjoint process which is determined by an infinite-time BSDE
+and introduce the adjoint process, which is determined by an infinite-time BSDE:
 $\mathrm{d}Y_{t}^{x,\beta}=-\left(\partial_{x}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})\right)\mathrm{d}t+Z_{t}^{x,\beta}\mathrm{d}B_{t}.$ (4.8)
-Applying Itô’s formula to the process $(e^{-rt}R_{t}Y_{t}^{x,\beta})$, by simple computation we can deduce that
+Applying Itô’s formula to the process $(e^{-rt}R_{t}Y_{t}^{x,\beta})$ and by simple computation, we can deduce that
 $\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}J(\beta+\epsilon\gamma)\right|_{\epsilon=0}=\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\partial_{a}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})\cdot\gamma_{t}\mathrm{d}t\right].$ (4.9)
@@
 $\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})=\min_{a\in A}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},a,Y_{t}^{x,\beta}).$ (4.10)
-Recalling our convexity assumptions on $b,f$ in Assumption (4.1), we know that the representative player’s optimal control $\beta$ takes a feedback form, that is
+Recalling our convexity assumptions on $b,f$ in Assumption 4.1, we know that the representative player’s optimal control $\beta$ takes a feedback form, that is,
 $\beta_{t}=\hat{\alpha}(X_{t}^{x,\beta},Y_{t}^{x,\beta}).$ (4.11)
-Now we consider the following two McKean–Vlasov FBSDEs:
+Now, we consider the following two McKean-Vlasov FBSDEs:
 $\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.12)
 $\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.13)
-Here (4.12) and (4.13) denote the state processes of social equilibrium and the representative player, respectively.
+Here, (4.12) and (4.13) denote the state processes of the social equilibrium and the representative player, respectively.
 Observe that their admissible controls both take the identical feedback form $\hat{\alpha}(x,y)$.
-We shall prove that when social equilibrium employ this feedback control, the representative player’s loss function is minimized if they use the same feedback form—thereby constituting a Nash equilibrium.
+We shall prove that when the social equilibrium employs this feedback control, the representative player’s loss function is minimized if they use the same feedback form, thereby constituting a Nash equilibrium.
 Theorem 4.2
-Let $(b,f)$ be differentiable in $(x,a)$, $\mathcal{H}$ be convex in $(x,a)$.
+Let $(b,f)$ be differentiable in $(x,a)$ and $\mathcal{H}$ be convex in $(x,a)$.
 Suppose that $\hat{\alpha}$ is Lipschitz continuous and that both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
-If we denote $\hat{\alpha}(X_{t}^{x},Y_{t}^{x})$ as $\alpha^{*}_{t}$ which is an admissible control in $\mathcal{A}$, then we have
+If we denote $\hat{\alpha}(X_{t}^{x},Y_{t}^{x})$ as $\alpha^{*}_{t}$, which is an admissible control in $\mathcal{A}$, then we have
 $J({\alpha}^{*})=\min_{\beta\in\mathcal{A}}J(\beta).$ (4.14)
@@
 $+r\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}(X_{t}^{x}-X_{t}^{x,\beta})\cdot Y_{t}^{x}\mathrm{d}t\right].$
-Since $X_{t}^{x},X_{t}^{x,\beta},Y_{t}^{x}$ are all belong to $L^{2}_{r}$, we can find a sequence of $T_{i}\rightarrow\infty$ such that
+Since $X_{t}^{x},X_{t}^{x,\beta},Y_{t}^{x}$ all belong to $L^{2}_{r}$, we can find a sequence of $T_{i}\rightarrow\infty$ such that
 $\mathbb{E}\left[e^{-rT_{i}}(X_{T_{i}}^{x}-X_{T_{i}}^{x,\beta})\cdot Y_{T_{i}}^{x}\right]\to 0.$ (4.17)
@@
 $J(\alpha^{*})-J(\beta)\leq 0$ (4.20)
-for all admissible control $\beta$. Thus we complete the proof.
+for all admissible controls $\beta$. Thus, we complete the proof.
 4.2 Solvability of mean field game FBSDEs
 In this subsection, we will find sufficient conditions for the existence and uniqueness of solutions to (4.12) and (4.13).
-Considering the linear case, we assume $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
-We require a technical lemma about the Lipschitz and convex property of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in ([1], Lemma 3.1).
+Considering the linear case, we assume that $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
+We require a technical lemma about the Lipschitz and convex properties of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in [1] (Lemma 3.1).
 Lemma 4.3
-Suppose $f_{1}$ is once continuously differentiable in $(x,a)$, $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$. And $f_{1}$ is $\eta$-convex in $a$, which means
+Suppose $f_{1}$ is once continuously differentiable in $(x,a)$, $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $f_{1}$ is $\eta$-convex in $a$, which means that
 $f_{1}(x,a^{\prime})-f_{1}(x,a)-(a^{\prime}-a)\cdot\partial_{a}f_{1}(x,a)\geq\eta|a^{\prime}-a|^{2},\quad~\mbox{for all $x\in\mathbb{R}$.}$ (4.21)
-Then it holds that
+Then, it holds that
 $|\hat{\alpha}(x,y)-\hat{\alpha}(x^{\prime},y^{\prime})|\leq\frac{\ell}{2\eta}|x^{\prime}-x|+\frac{|b_{3}|}{2\eta}|y^{\prime}-y|,$ (4.22)
-and for some $a_{0}\in A$
+and for some $a_{0}\in A$,
 $|\hat{\alpha}(x,y)|\leq\eta^{-1}(|\partial_{a}f_{1}(x,a_{0})|+|b_{2}y|)+|a_{0}|.$ (4.23)
-Furthermore, if $A=\mathbb{R}$ and $\partial_{a}f$ is $\zeta$-Lipchitz in $a$, it follows that
+Furthermore, if $A=\mathbb{R}$ and $\partial_{a}f$ is $\zeta$-Lipschitz in $a$, it follows that
 $b_{3}(y^{\prime}-y)\cdot\left(\hat{\alpha}(x,y^{\prime})-\hat{\alpha}(x,y)\right)\leq-\frac{2b_{3}^{2}\eta}{\zeta^{2}}(y^{\prime}-y)^{2}.$ (4.24)
@@
 Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$.
-Under the following conditions, Assumption 3.1 and 4.1 are satisfied, and both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
+Under the following conditions, Assumptions 3.1 and 4.1 are satisfied and both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
 (i) There exists a positive constant $k$ such that $|b_{2}|\leq k$ and $b_{1}\geq k-\frac{r}{2}$.
@@
 $f_{1}(x,a)$ is once continuously differentiable and of at most quadratic growth in $(x,a)$.
-(ii) $f_{0}(x,\mu)$ is convex in $x$ and there exists a positive constant $b_{4}>0$, such that
-$|\partial_{x}f_{0}(x,\mu)-\partial_{x}f_{0}(x^{\prime},\mu^{\prime})|<b_{4}\left(|x-x^{\prime}|+W_{2}(\mu,\mu^{\prime})\right).$ (4.25)
-(iii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds
+(ii) $f_{0}(x,\mu)$ is convex in $x$, and there exists a positive constant $b_{4}>0$ such that
+$|\partial_{x}f_{0}(x,\mu)-\partial_{x}f_{0}(x^{\prime},\mu^{\prime})|<b_{4}\left(|x-x^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})\right).$ (4.25)
+(iii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds:
 $f_{1}(x^{\prime},a^{\prime})-f_{1}(x,a)-\partial_{(x,a)}f_{1}(x,a)\cdot(x^{\prime}-x,a^{\prime}-a)$ (4.26)
 $\geq$ $\iota(x^{\prime}-x)^{2}+\eta(a^{\prime}-a)^{2}.$
-(iv) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipchitz continuous in $(x,a)$.
+(iv) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipschitz continuous in $(x,a)$.
 (v) $A=\mathbb{R}$, and it holds that
@@
 Remark 4.5
-If we set $f_{0}(x,\mu)=b_{4}x\bar{\mu},f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C,\ell=0$. Then the requirement in (4.27) becomes
+If we set $f_{0}(x,\mu)=b_{4}x\bar{\mu},f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C,\ell=0$. Then, the requirement in (4.27) becomes
 $\min\left\{2A-\frac{|b_{2}|}{2}-|b_{4}|,\frac{b_{3}^{2}}{2C}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.28)
@@
 Proof.
-It’s clear that Assumption 3.1 is satisfied.
-And by Lemma 4.3, $\hat{\alpha}$ is Lipchitz.
-In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+f_{0}(x,\mu)+f_{1}(x,a)$ is convex in $(x,a)$ according to assumption on $f_{0},f_{1}$, thus Assumption 4.1 is satisfied.
-To prove that the BSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
-We consider the FBSDE (4.12), and set
+It is clear that Assumption 3.1 is satisfied.
+And by Lemma 4.3, $\hat{\alpha}$ is Lipschitz.
+In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+f_{0}(x,\mu)+f_{1}(x,a)$ is convex in $(x,a)$ according to the assumption on $f_{0},f_{1}$; thus, Assumption 4.1 is satisfied.
+To prove that the FBSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
+We consider the FBSDE (4.12) and set
 $B(t,x,y,\mu)=b_{1}x+b_{2}\overline{\mu}+b_{3}\hat{\alpha}(x,y),$ (4.29)
 $F(t,x,y,\mu)=b_{1}y+\partial_{x}f_{0}(x,\mu)+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
-Take four arbitrary square integrable random variables $X,Y,X^{\prime},Y^{\prime}$.
+Take four arbitrary square-integrable random variables $X,Y,X^{\prime},Y^{\prime}$.
 Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ .
@@
 $<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
+By Lemma 2.2, we know both FBSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
 5 Master equation representation
@@
 $\partial_{y}H(x,\mu,y)=b(x,\mu,\hat{\alpha}(x,y)).$ (5.2)
-Assume the master equation (1.7) has a classical solution $U(x,\mu)\in{\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ with $H$ defined above, and $\hat{b}(x,\mu)=b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$.
+Assume that the master equation (1.8) has a classical solution $U(x,\mu)\in C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ with $H$ defined above such that $\partial_{x}U(x,\mu)\in C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ satisfying PDE (1.12) and $\hat{b}(x,\mu)=b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$.
 For any $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$, suppose the following SDE admits a unique solution in $L^{2}_{r}$:
 $\mathcal{X}_{t}^{\xi}=\xi+\int_{0}^{t}b\left(\mathcal{X}_{s}^{\xi},\mathcal{L}_{\mathcal{X}_{s}^{\xi}},\hat{\alpha}\left(\mathcal{X}_{s}^{\xi},\partial_{x}U\left(\mathcal{X}_{s}^{\xi},\mathcal{L}_{\mathcal{X}_{s}^{\xi}}\right)\right)\right)\mathrm{d}s+B_{t}.$ (5.3)
-Then we take $\mathcal{Y}_{t}^{\xi}=\partial_{x}U(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}})$.
+Then, we take $\mathcal{Y}_{t}^{\xi}=\partial_{x}U(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}})$.
 By differentiating both sides of the master equation (1.8) with respect to $x$ and applying Itô’s formula, we obtain
@@
 Theorem 5.1
-Assume the master equation (1.7) admits a classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
+Assume that the master equation (1.8) admits a classical solution $U(x,\mu)\in C^{2,1}(\mathbb{R}\times{\cal P}_{2})$, which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$
 $\begin{cases}X_{t}^{\xi}=\xi+\int_{0}^{t}b(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}},\hat{\alpha}(X_{s}^{\xi},\partial_{x}U(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}})))\mathrm{d}t+B_{t},\\ X_{t}^{x}=x+\int_{0}^{t}b(X_{s}^{x},\mathcal{L}_{X_{s}^{x}},\hat{\alpha}(X_{s}^{x},\partial_{x}U(X_{s}^{x},\mathcal{L}_{X_{s}^{x}})))\mathrm{d}t+B_{t}.\end{cases}$ (5.8)
-Then for every initial distribution $\mu_{0}$, $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium, and $U(x,\mu_{0})$ is exactly the value function given $\alpha^{*}$.
-Proof.  Given $\alpha^{*}$ and initial $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu_{0})$, the states of social equilibrium is governed by the following SDE:
+Then, for every initial distribution $\mu_{0}$, $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium, and $U(x,\mu_{0})$ is exactly the value function given $\alpha^{*}$.
+Proof.  Given $\alpha^{*}$ and initial $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu_{0})$, the state of social equilibrium is governed by the following SDE:
 $X_{t}^{\xi}=\xi+\int_{0}^{t}\partial_{y}H(X_{s}^{\xi},\rho_{s},\partial_{x}U(X_{s}^{\xi},\rho_{s}))\mathrm{d}s+B_{t},\quad\quad\rho_{s}\triangleq\mathcal{L}_{X_{s}^{\xi}}.$ (5.9)
@@
 for every feasible control $\beta$, and the equality holds when $\beta_{t}=\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t}))$.
-This shows $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is an Nash equilibrium.
-And we have
-$U(x,\mu_{0})=\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t})))\mathrm{d}t,$ (5.13)
-this completes the proofs of our desired results.
-At the end of this section, we provide an example of a solvable elliptic master equation.
+This shows that $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium.
+In addition, we have
+$U(x,\mu_{0})=\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t})))\mathrm{d}t.$ (5.13)
+This completes the proof of our desired results.
+At the end of this section, we provide an example of a solvable elliptic master equation and demonstrate that its solutions are generally nonunique.
 Set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a,f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, where $b_{1},b_{2},b_{3},b_{4},A,C$ are constants satisfying all conditions in Theorem 4.4 and Remark 4.5.
-Then we have $\hat{\alpha}(x,y)=-\frac{b_{3}y}{2C}$ and
+Then, we have $\hat{\alpha}(x,y)=-\frac{b_{3}y}{2C}$ and
 $H(x,\mu,y)=(b_{1}x+b_{2}\bar{\mu})\cdot y+b_{4}x\bar{\mu}+Ax^{2}-\frac{b_{3}^{2}}{4C}y^{2}.$ (5.14)
@@
 $\begin{split}rU(x,\mu)=&(b_{1}x+b_{2}\bar{\mu})\cdot\partial_{x}U(x,\mu)+b_{4}x\bar{\mu}+Ax^{2}\\ &-\frac{b_{3}^{2}}{4C}(\partial_{x}U(x,\mu))^{2}+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})(b_{1}\tilde{\xi}+b_{2}\bar{\mu}-\frac{b^{2}_{3}}{2C}\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (5.15)
-Theorem 5.2
-The master equation (5.15) admits a classical solution $U(x,\mu)$ in ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$.
-Proof.  We assume that the solution takes the form
+We assume that the solution takes the form
 $U(x,\mu)=a_{1}x^{2}+a_{2}x\bar{\mu}+a_{3}(\bar{\mu})^{2}+a_{4},$ (5.16)
-where $a_{1},a_{2},a_{3},a_{4}$ are constants remain to be determined. Then we have
+where $a_{1},a_{2},a_{3},a_{4}$ are constants that remain to be determined. Then we have
 $\begin{split}&\partial_{x}U(x,\mu)=2a_{1}x+a_{2}\bar{\mu},\quad\quad\quad\partial_{xx}U(x,\mu)=2a_{1},\\ &\partial_{\mu}U(x,\mu,\tilde{x})=a_{2}x+2a_{3}\bar{\mu},\quad\quad\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{x})=0.\end{split}$ (5.17)
@@
 $\begin{cases}ra_{1}=2b_{1}a_{1}+A-\frac{b_{3}^{2}}{C}a_{1}^{2},\\ ra_{2}=2b_{2}a_{1}+b_{1}a_{2}+b_{4}-\frac{b_{3}^{2}}{C}a_{1}a_{2}+a_{2}(b_{1}+b_{2}-\frac{b_{3}^{2}}{C}a_{1}-\frac{b_{3}^{2}}{2C}a_{2}),\\ ra_{3}=b_{2}a_{2}-\frac{b_{3}^{2}}{4C}a_{2}^{2}+2a_{3}(b_{1}+b_{2}-\frac{b_{3}^{2}}{C}a_{1}-\frac{b_{3}^{2}}{2C}a_{2}),\\ ra_{4}=a_{1}.\end{cases}$ (5.19)
-Then, we can solve for $a_{1},a_{2},a_{3},a_{4}$.
-Remark 5.3
-We have
-$\hat{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))=(b_{1}-\frac{a_{1}b_{3}^{2}}{C})x+(b_{2}-\frac{a_{2}b_{3}^{2}}{2C})\bar{\mu}.$ (5.20)
-Since the SDE (5.3) now becomes linear, if we further require $a_{1}>0,a_{2}>0$, it’s not hard to see that the SDE admits a unique solution in $L_{r}^{2}$.
+Note that $a_{1},a_{2},a_{3}$, and $a_{4}$ can be solved sequentially, and the equations for $a_{1}$ and $a_{2}$ are quadratic.
+Since $A>0$ and $\frac{b_{3}^{2}}{C}<0$, the equation for $a_{1}$ must have one positive and one negative root.
+If the coefficients are sufficiently well-behaved, for instance, $2b_{2}a_{1}+b_{4}>0$, then the equation for $a_{2}$ will also have two distinct real roots.
+Therefore, the system of equations (5.19) can have at most four solutions.
+It must be emphasized that there exists at most one set of solutions for which the solution of SDE (5.3) lies in $L_{r}^{2}$, as the $L_{r}^{2}$ solution of FBSDE (4.12) is unique.
+In a companion article, we establish the uniqueness of solutions to the master equation under certain monotonicity and growth assumptions.
+Example 5.2
+We consider a case that $r=2,b_{1}=b_{2}=b_{4}=0,b_{3}=2,A=2,C=1$, then equation (5.15) becomes
+$2U(x,\mu)=$ $2x^{2}-\left(\partial_{x}U(x,\mu)\right)^{2}+\frac{1}{2}\partial_{xx}U(x,\mu)$ (5.20)
+$+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})-2\partial_{\mu}U(x,\mu,\tilde{\xi})\cdot\partial_{x}U(\tilde{\xi},\mu)\right]$
+Solving the corresponding Equation (5.19) yields the following solutions:
+$U_{1}(x,\mu)=\frac{1}{2}x^{2}+\frac{1}{4},$ (5.21)
+$U_{2}(x,\mu)=\frac{1}{2}x^{2}-3x\bar{\mu}+\frac{3}{2}(\bar{\mu})^{2}+\frac{1}{4},$
+$U_{3}(x,\mu)=-x^{2}-\frac{1}{2},$
+$U_{4}(x,\mu)=-x^{2}+3x\bar{\mu}-\frac{3}{2}(\bar{\mu})^{2}-\frac{1}{2}.$
+However, to ensure that the solution of Equation (5.3) , now written as
+$\mathrm{d}X_{t}=-4\partial_{x}U(X_{t},{\cal L}_{X_{t}})\mathrm{d}t+\mathrm{d}B_{t},\quad X_{0}=\xi,$ (5.22)
+belong to $L_{2}^{2}$, only the solution $U_{1}(x,\mu)=\frac{1}{2}x^{2}+\frac{1}{4}$ satisfies this condition.
 6 Viscosity solution to distribution-dependent elliptic PDE
@@
 $\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (6.2)
-Here $H(x,\mu,y)$ is defined in (5.1) and the above equations are the same as (1.5) and (1.6).
-We define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ and attempt to prove that it satisfy the following elliptic PDE:
+Here, $H(x,\mu,y)$ is defined in (5.1), and the above equations are the same as (1.6) and (1.7).
+We define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ and attempt to prove that it satisfies the following elliptic PDE in the viscosity sense:
 $\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (6.3)
 This equation is derived by taking the partial derivative of both sides of the master equation (1.8) with respect to $x$. We have the relationship $\mathcal{U}(x,\mu)=\partial_{x}U(x,\mu)$.
+Now, let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
+$(\mathcal{L}\mathcal{U})[\mathcal{U}](x,\mu)+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.4)
+where
+$\begin{split}(\mathcal{L}\Phi)[\Psi](x,\mu)\triangleq&\partial_{y}H(x,\mu,\Psi(x,\mu))\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.5)
+and
+$F(x,\mu,\Psi(x,\mu))\triangleq\partial_{x}H(x,\mu,\Psi(x,\mu))-r\Psi(x,\mu).$ (6.6)
+Definition 6.1
+A function $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ is said to be a test function if the quantities:
+[…]
+and
+[…]
+are finite, uniformly in $(x,\mu)$ in any compact subset of $\mathbb{R}\times\mathcal{P}_{2}$.
+Definition 6.2
+Let $\mathcal{U}\in C(\mathbb{R}\times\mathcal{P}_{2})$.
+Then $\mathcal{U}$ is called a viscosity subsolution (resp.
+supersolution) of PDE (6.4) if, whenever $\Psi$ is a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp.
+minimum) of $\mathcal{U}-\Psi$, we have
+$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.9)
+(respectively,
+$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.10)
+The function $\mathcal{U}$ is called a viscosity solution of PDE (6.4) if it is both a viscosity subsolution and a viscosity supersolution.
 The proof of $\mathcal{V}$ being a viscosity solution to equation (6.3) rests on two fundamental results:
@@
   • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$.
+We will next establish them sequentially in this section and ultimately prove that $\mathcal{V}$ is a viscosity solution to equation (6.3)
 6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
@@
 $\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.11)
-We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}\in\mathbb{L}^{2}({\cal F}_{0})$.
+We know that under Assumption 2.1, the above FBSDE has a unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}\in\mathbb{L}^{2}({\cal F}_{0})$.
 We then give the definitions of strong and weak uniqueness.
 Definition 6.3 (Strong uniqueness)
-We say that the strong uniqueness holds for FBSDE (6.4) if on any admissible set-up $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
+We say that the strong uniqueness holds for FBSDE (6.11) if on any admissible setup $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
 $(X^{1}_{t},Y^{1}_{t},Z^{1}_{t})_{t\geq 0},\quad(X^{2}_{t},Y^{2}_{t},Z^{2}_{t})_{t\geq 0}$ (6.12)
@@
 For any two set-ups $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1},\mathbb{F}^{1})$ and $(\Omega^{2},\mathcal{F}^{2},\mathbb{P}^{2},\mathbb{F}^{2})$ with inputs $(X_{0}^{1},B^{1})$ and $(X_{0}^{2},B^{2})$, $X_{0}^{1},X_{0}^{2}$ have the same law on $\mathbb{R}$.
-We say the weak uniqueness holds for FBSDE (6.4) if the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on corresponding set-ups have the same distribution.
-We use the same scheme as the one developed by Yamada and Watanabe to prove that path-wise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
+We say the weak uniqueness holds for FBSDE (6.11) if for the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},Z^{1}_{t})_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},Z^{2}_{t})_{t\geq 0}$ on corresponding set-ups, the processes $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ have the same distribution.
+We use the same scheme as the one developed by Yamada and Watanabe to prove that pathwise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
 Theorem 6.5
 Assume that on $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B_{t})_{t\geq 0}$, the FBSDE has a unique strong solution $(X_{t},Y_{t},Z_{t})$.
-Then the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
-Proof.  At first we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
+Then, the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
+Proof.  First, we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
 $d(\omega^{1},\omega^{2})=\sum_{n\geq 0}2^{-n}\sup_{t\in[0,n]}\max(|\omega_{t}^{1}-\omega_{t}^{2}|,1).$ (6.14)
@@
     $Q^{i}(G\times F)=\int_{G}q^{i}(\omega_{\text{input}};F)Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}).$ (6.17)
-Next, we need a enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
+Next, we need an enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
 $\Omega_{\text{total}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}\times\Omega_{\text{output}},$ (6.18)
@@
 the last inequality following from Fubini’s theorem.
-Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augumentation under $Q$ of the canonical filtration
+Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augmentation under $Q$ of the canonical filtration
 ${\cal H}_{t}=\{(\xi,w_{s},x^{1}_{s},y^{1}_{s},\zeta^{1}_{s},x^{2}_{s},y^{2}_{s},\zeta^{2}_{s});0\leq s\leq t\}$ (6.23)
 on $\Omega_{\text{total}}$.
-It’s easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable.
+It is easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable.
 Moreover, for $i\in\{1,2\}$,
 $Q\left\{\omega\in\Omega_{\text{total}};(x,w,x^{i},y^{i},\zeta^{i})\in A\right\}=Q^{i}\left\{(X_{0}^{i},B^{i},X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\right\};\quad A\in{\cal B}(\Omega_{\text{canon}}).$ (6.24)
-Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in ([6], Remark 1.6). Let us firstly define
+Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in [6] (Remark 1.6). Let us firstly define
 $\pi_{t}:C([0,\infty);\mathbb{R})\rightarrow C([0,t];\mathbb{R}),\quad h\rightarrow h|_{[0,t]},$ (6.25)
@@
 Noting that ${\cal H}^{t}={\cal B}(\mathbb{R})\otimes{\cal K}_{t}\otimes{\cal K}^{\prime}_{t}\otimes{\cal K}^{\prime}_{t}$, we conclude that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion.
-At last, applying the same procedure on $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in ([5], Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
+At last, applying the same procedure to $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in [5] (Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
 $\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}G(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.33)
@@
 We can say that $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$.
-6.2 Regularity of the FBSDE solutions
+6.2 Connection with distribution-dependent elliptic PDE
+For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$. We first present the continuity result for $\mathcal{V}$.
 Lemma 6.7
-For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$.
-Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
+Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$ such that
 $|\mathcal{V}(x,\mu)-\mathcal{V}(x^{\prime},\mu^{\prime})\leq C(|x-x^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime}))$ (6.34)
@@
 $\begin{split}(\hat{Y}^{x,\xi}_{0})^{2}\leq&C_{1}C_{2}\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}+(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\\ \leq&\frac{1}{2}(\hat{Y}^{x,\xi}_{0})^{2}+\frac{C_{1}^{2}C_{2}^{2}}{2}(\hat{X}^{x,\xi}_{0})^{2}+(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t.\end{split}$ (6.38)
-Thus
+Thus,
 $(\hat{Y}^{x,\xi}_{0})^{2}\leq C_{1}^{2}C_{2}^{2}(\hat{X}^{x,\xi}_{0})^{2}+2(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t.$ (6.39)
@@
 $\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\leq C_{3}\mathbb{E}|\xi_{1}-\xi_{2}|^{2}.$ (6.40)
-So we can deduce that
+So, we can deduce that
 $|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathbb{E}|\xi_{1}-\xi_{2}|^{2}\right).$ (6.41)
 Since $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$, we have
-$|\mathcal{V}(x_{1},\mu_{1})-\mathcal{V}(x_{2},\mu_{2})|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathcal{W}_{2}^{2}(\mu_{1},\mu_{2})\right),$ (6.35)
-then we get the desired result.
-6.3 Connection with distribution-dependent elliptic PDE
-Now let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
-$(\mathcal{L}\mathcal{U})[\mathcal{U}](x,\mu)+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.36)
-where
-$\begin{split}(\mathcal{L}\Phi)[\Psi](x,\mu)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
-and
-$F(x,\mu,\Psi(x,\mu))\triangleq\partial_{x}H(x,\mu,\Psi(x,\mu))-r\Psi(x,\mu).$ (6.38)
-Definition 6.6
-A function $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ is said to be a test function if the quantities:
-[…]
-and
-[…]
-are finite, uniformly in $(x,\mu)$ in any compact subset of $\mathbb{R}\times\mathcal{P}_{2}$.
-Definition 6.7
-Let $\mathcal{U}\in C(\mathbb{R}\times\mathcal{P}_{2})$.
-Then $\mathcal{U}$ is called a viscosity subsolution (resp.
-supersolution) of PDE (6.36) if, whenever $\Psi$ is a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp.
-minimum) of $\mathcal{U}-\Psi$, we have
-$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.41)
-(respectively
-$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.42)
-The function $\mathcal{U}$ is called a viscosity solution of PDE (6.36) if it is both a viscosity subsolution and a viscosity supersolution.
-We now assert that
+$|\mathcal{V}(x_{1},\mu_{1})-\mathcal{V}(x_{2},\mu_{2})|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathcal{W}_{2}^{2}(\mu_{1},\mu_{2})\right).$ (6.42)
+Then, we get the desired result.
+We now assert the following theorem.
 Theorem 6.8
-Assume that both FBSDEs (6.1) and (6.2) admit unique solutions in $L^{2}_{r}$, $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$.
-Then we define the function $\mathcal{V}(x,\mu)\triangleq Y_{0}^{x,\xi}$ with $\mathcal{L}_{\xi}=\mu$, and assume $\mathcal{V}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$.
-The function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.36).
-Proof.  We only show that $\mathcal{V}$ is a viscosity subsolution of PDE (6.36). A similar argument will show that it is also a viscosity supersolution of (6.36).
-Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$.
-Let $\Psi$ be a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $\mathcal{U}-\Psi$.
+Considering all assumptions in Theorem 4.4 are satisfied, the function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.4).
+Proof.  We only show that $\mathcal{V}$ is a viscosity subsolution of PDE (6.4). A similar argument will show that it is also a viscosity supersolution of PDE (6.4).
+Due to the uniqueness of solutions of FBSDEs, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$.
+Let $\Psi$ be a test function and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $\mathcal{U}-\Psi$.
 We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$.
@@
 $(\mathcal{L}\Psi)[\mathcal{V}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.43)
-It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
+It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$ such that for all $(x,\mu)\in O$,
 $\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)[\mathcal{V}](x,\mu)+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.44)
-Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu^{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2).
+Taking an initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu^{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$, which are solutions to FBSDEs (6.1) and (6.2), respectively.
 For some $T>0$, let $\tau$ denote the stopping time
@@
 $(\bar{Y}\left(t\right),\bar{Z}\left(t\right))\triangleq(Y^{x^{0},\xi^{0}}_{t\wedge\tau},I_{\left[0,\tau\right]}(t)Z^{x^{0},\xi^{0}}_{t}),\quad 0\leq t\leq T,$ (6.46)
-is the solution of BSDE
+is the solution of the BSDE
 $\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\rho_{s},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\rho_{s}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.47)
@@
 $(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\rho_{t\wedge\tau}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\rho_{t\wedge\tau}))$ (6.48)
-is the solution of BSDE
+is the solution of the BSDE
 $\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi[\mathcal{V}](X^{x^{0},\xi^{0}}_{s},\rho_{s})\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.49)
@@
 $\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.52)
-Therefore
+Therefore,
 $\tilde{Y}_{0}=\mathbb{E}\left[\tilde{Y}_{\tau}+\int_{0}^{\tau}\beta_{s}\mathrm{d}s\right]$ (6.53)
-Now from the choice of $O$ and $\tau$, a.s.
+Now, from the choice of $O$ and $\tau$, a.s.
 $\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.54)
 Consequently, $\tilde{Y}_{0}=\Psi(x^{0},\mu^{0})-\mathcal{V}(x^{0},\mu^{0})>0$, which contradicts the earlier assumption.
+Acknowledgements
+Song Y. is financially supported by National Key R&D Program of China (No. 2024YFA1013503 & No. 2020YFA0712700) and the National Natural Science Foundation of China (No. 12431017).